### v4

# Almost mathematics of pointed symmetric monoidal model categories by Smith ideal theory

Abstract.
This article is a generalization of a result in Quillen’s note [Qui96] “Module theory over non-unital rings” giving a one-to-one correspondence between bilocalization of abelian categories of modules and idempotent ideals of the base ring. Faltings [Fal88]; Gabber and Ramero [GR03] established almost mathematics, the same as Quillen’s bilocalization of a category of modules by nil-modules.
In this paper, by using the theory of Smith ideals mentioned in Hovey [Hov14], we consider almost mathematics of symmetric monoidal pointed model categories and prove a weak analogue of the one-to-one correspondence in Quillen’s note [Qui96].
Key words and phrases:
Symmetric monoidal model categories, Smith ideals, almost mathematics, Bousfield localization and colocalization
1. Introduction
Faltings [Fal88] first introduced one of the topics of algebra theory almost mathematics. Later, Gabber and Ramero provided a textbook [GR03] of detailed foundation of almost mathematics, where they called those materials of the theory almost ring theory, almost modules, almost algebras, and almost homotopy algebras. Referring Gabber and Ramero’s book, almost mathematics is obtained by a localization of an abelian category of modules by almost isomorphisms: Let $V$ be a unital commutative ring with an idempotent ideal $\mathcal{m}$ of $V$. A $V$-module $M$ is said to be almost zero if $M$ is killed by $\mathcal{m}$, and almost mathematics is a theory of algebra working by localizing the category of $V$-modules by the Serre subcategory of almost zero modules. Almost isomorphisms are morphisms whose kernels and cokernels are both almost zero. In the localized category, module objects are called almost modules and algebra objects are called almost algebras.
Independently, Quillen considered the same theory of almost mathematics in unpublished note [Qui96] as linear algebra over non-unital rings, which he characterized as bilocalization - both localization and colocalization - of abelian categories of modules. Further, Quillen proved the following one-one correspondence between idempotent ideals and reflexive bilocalization functors.
Theorem 1.1.
([Qui96] Proposition 6.5) Let $V$ be a commutative ring with a multiplicative unit. There is a one-to-one correspondence between Serre subcategories $\mathcal{S}$ of $\mathrm{Mod}_{V}$ that the localization $F:\mathrm{Mod}_{V}\to\mathrm{Mod}_{V}/\mathcal{S}$ is also a colocalization, and idempotent ideals of $V$. $\Box$
The purpose of this article is to establish an almost mathematics theory of symmetric monoidal model categories, where we prove a monoidal categorical analogue of Theorem 1.1. Namely, by utilizing the commutative Smith ideal theory developed by Hovey [Hov14] and White–Yau [WY24], we introduce homotopy analogues of idempotent ideals and almost isomorphisms, which we refer to as homotopically idempotent Smith ideals and almost weak equivalences. The main results of this paper are as follows, as mentioned in Theorem 2.11.
Theorem A.
Let $\mathcal{M}$ be a pointed symmetric monoidal model category with a monoidal unit object $V$, $j:I\to V$ be a homotopically idempotent commutative Smith ideal. Then the internal-hom functor
$\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-):\mathcal{M}\to\mathcal{M}$
is equivalent to the Bousfield bilocalization functor by almost weak equivalences.
Since $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ is a right Quillen functor, this theorem shows us that the Bousfield localization by almost weak equivalences is both a left and right Quillen functor.
Conversely, in the case that the monoidal unit generates the stable symmetric monoidal model category in the sense of MacLane [Mac88, Chapter V, Section 7, p.127], we can construct a homotopically idempotent ideal (Proposition 4.5 and Theorem 4.7).
Furthermore, we generalize some result of almost algebras for model category setting by using Smith ideal language. Again, let $V$ be a commutative ring with an idempotent ideal $\mathfrak{m}$. We assume that $\tilde{\mathfrak{m}}$ is $V$-flat. Let $\mathrm{CAlg}(\mathrm{Mod}_{V})$ denote the category of $V$-algebras and $\mathrm{CAlg}(\mathrm{Mod}_{V}^{\rm al})$ the category of almost algebras. Gabber and Ramero defined the functor
$(-)_{!!}:\mathrm{CAlg}(\mathrm{Mod}_{V})\to\mathrm{CAlg}(\mathrm{Mod}_{V}^{\rm al})$
by setting
$A_{!!}=V\amalg_{\tilde{\mathfrak{m}}}(\tilde{\mathfrak{m}}\otimes_{V}A)$
for any almost unital $V$-algebra $A$ in [GR03, Section 2.2]. Those morphisms $\tilde{\mathfrak{m}}\to V\to A$ and $\tilde{\mathfrak{m}}\to\tilde{\mathfrak{m}}\otimes_{V}A\to A$ induce a ring homomorphism $A_{!!}\to A$, which is almost isomorphism. Due to [GR03, Proposition 2.2.29], the almost unitalization $(-)_{!!}$ is a left adjoint of the localization $\mathrm{CAlg}(\mathrm{Mod}_{V})\to\mathrm{CAlg}(\mathrm{Mod}_{V}^{\rm al})$. We call the left adjoint $(-)_{!!}$ almost unitalization in this paper. By using Smith ideal theory, we generalize the almost unitalization functor $(-)_{!!}$ that is right adjoint of the localization functor.
In Section 2, we introduce the theory of almost mathematics of pointed symmetric monoidal model categories; homotopically idempotent commutative Smith ideals, homotopically almost zero objects, and homotopically almost weak equivalences.
In Section 3, we generalize the above almost unitalization functor by Smith ideal theory language, and we prove that it is a left adjoint of the Bousfield localization functor by the class of almost weak equivalences defined in Section 2, as Theorem 3.6 and Corollary 3.7.
In Section 4, as the inverse direction of Theorem 2.11, we demonstrate constructing homotopically idempotent Smith ideal from given symmetric monoidal bilocalization functor on symmetric monoidal model categories under some mild conditions, and the result is proven as Theorem 4.7.
Finally, in Section 5, we remark that the cokernel-and-kernel functors in the (operadic) Smith ideal theory induces a categorical equivalence between non-unital algebra objects and augmented algebra objects of stable symmetric monoidal model categories.
1.1. Example: The almost mathematics of a stable model category of chain complexes
In this section, we consider the model category of chain complexes of modules as an example of almost mathematics of model categories. We fix a commutative base ring $V$ and an idempotent ideal $\mathfrak{m}$ ($\mathfrak{m}^{2}=\mathfrak{m}$) of $V$, and we assume that the tensor product $\tilde{\mathfrak{m}}=\mathfrak{m}\otimes_{V}\mathfrak{m}$ is a flat $V$-module. Let $\operatorname{Ch}(V)$ denote the category of chain complexes of $V$-modules. Then $\operatorname{Ch}(V)$ has a canonical model structure called the injective model structure that cofibrations are degree-wise injections and weak equivalences are quasi-isomorphisms. Since we assume that $\tilde{\mathfrak{m}}$ is a flat $V$-module, the functor $\tilde{\mathfrak{m}}\otimes_{V}(-):\operatorname{Ch}(V)\to\operatorname{Ch}(V)$ is exact and a left Quillen functor, whose right adjoint is the mapping complex functor $\mathrm{Map}_{V}(\tilde{\mathfrak{m}},\,-)$.
Example 1.2.
A perfectoid commutative valuation ring is defined to be a valuation ring of a perfectoid field defined in Scholze’s paper [Sch12, p.246, Definition 1.2]. Let $V$ be a perfectoid valuation ring. Then the set of topologically nilpotent elements $\mathfrak{m}$ of $V$ is an idempotent maximal ideal and a flat $V$-module, explained in, for example, Bhatt’s lecture note [Bha17, p.18, Example 4.1.3]. In general, the product map $\tilde{\mathfrak{m}}\to\mathfrak{m}$ is an isomorphism if $\mathfrak{m}$ is a flat $V$-module.
In the category $\mathrm{Ch}(V)$ of almost weak equivalences is defined as the following:
Definition 1.3.
A morphism $f:M\to N$ of $V$-complexes is said to be an almost quasi-isomorphism if the induced morphism $\tilde{\mathfrak{m}}\otimes_{V}f:\tilde{\mathfrak{m}}\otimes_{V}M\to\tilde{\mathfrak{m}}\otimes_{V}N$ is a quasi-isomorphism. Let $\operatorname{Ch}(V)^{\rm al}$ denote the Bousfield localization of $\operatorname{Ch}(V)$ by the class of almost quasi-isomorphisms. A chain complex $E$ is said to be almost local if $E$ is a local object with respect to almost quasi-isomorphisms. That is, for any almost quasi-isomorphism $f:M\to N$, the induced map
$f^{*}:\mathrm{Hom}_{D(V)}(N,\,E)\to\mathrm{Hom}_{D(V)}(M,\,E)$
is an isomorphism in the derived category $D(V)$ of $\operatorname{Ch}(V)$.
In Quillen’s note [Qui96], a $V$-module $E$ that the induced homomorphism $E\to\mathrm{Hom}_{V}(\mathfrak{m},\,E)$ is an isomorphism is said to be closed, and closed modules are almost local objects and vice versa. The following lemma characterizes almost local objects.
Proposition 1.4.
Let $E$ be a chain complex of $V$-modules. Then $E$ is almost local if and only if the product $\mu_{E}:\tilde{\mathfrak{m}}\otimes_{V}E\to E$ induces a quasi-isomorphism
$E\to\mathrm{Map}_{V}(\tilde{\mathfrak{m}},\,E).$
  • proof.
    The only if direction is clear by definition of almost local objects.
Conversely, assume that $E$ is almost local. Since the canonical morphism $\tilde{\mathfrak{m}}\otimes_{V}M\to M$ is an almost quasi-isomorphism for any $M\in\operatorname{Ch}(V)$, then the induced morphism
$\mathrm{Hom}_{D(V)}(M,\,E)\to\mathrm{Hom}_{D(V)}(M,\,\mathrm{Map}_{V}(\tilde{\mathfrak{m}},\,E))\simeq\mathrm{Hom}_{D(V)}(\tilde{\mathfrak{m}}\otimes_{V}M,\,E).$
is an isomorphism. By the Yoneda lemma, $E\to\mathrm{Map}_{V}(\tilde{\mathfrak{m}},\,E)$ is a quasi-isomorphism. $\Box$
Theorem 1.5.
Then the mapping complex functor
$\mathrm{Map}_{V}(\tilde{\mathfrak{m}},\,-):\operatorname{Ch}(V)\to\operatorname{Ch}(V)^{\rm al}$
is equivalent to the localization functor $L^{\rm al}:\operatorname{Ch}(V)\to\operatorname{Ch}(V)^{\rm al}$ on the derived categories.
  • proof.
    This is the special case of Theorem 2.11. $\Box$
Acknowledgments
I would like to thank anonymous referees for giving me innovative ideas and constructive comments to use Hovey’s Smith ideal theory and White and Yau’s operadic Smith ideal theory.
2. Almost mathematics of pointed symmetric monoidal model categories
In this section, referring White and Yau’s work [WY24], we recall definition and some properties of symmetric monoidal model model categories and (commutative) Smith ideals of them. Next, in order to establish almost mathematics, we define homotopically idempotent Smith ideals, and we prove this paper’s main result Theorem 2.11.
[…]
Let $\mathcal{M}$ be a model category. For any object $M\in\mathcal{M}$, let $Q(M)\to M$ denote a cofibrant replacement and $M\to R(M)$ a fibrant replacement. It is well-known that the following adjunctions
${\mathcal{M}^{\rm cof}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R}$$\scriptstyle{Q}$${\mathcal{M}^{\rm fib}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
are Quillen equivalences, where $\mathcal{M}^{\rm cof}$ (resp. $\mathcal{M}^{\rm fib}$) denotes the full subcategory spanned by all cofibrant (resp. fibrant) objects.
2.1. Commutative Smith ideals of symmetric monoidal model categories
A symmetric monoidal model category $\mathcal{M}$ is a model category with a symmetric monoidal structure $-\otimes-:\mathcal{M}\times\mathcal{M}\to\mathcal{M}$ satisfying the following two axioms:
  1. (1)
    Unit axiom: The monoidal unit $V$ of $\mathcal{M}$ satisfies the condition that, for any cofibrant object $M$ and cofibrant replacement of $Q(V)\to V$, the induced morphisms $Q(V)\otimes M\to V\otimes M\to M$ and $M\otimes Q(V)\to M\otimes V\to M$ are weak equivalences.
  2. (2)
    Pushout-product axiom: For any cofibrations $f:X_{0}\to X_{1}$ and $g:Y_{0}\to Y_{1}$, the pushout product $f\Box g:(X_{0}\otimes Y_{1})\amalg_{X_{0}\otimes Y_{0}}(X_{1}\otimes Y_{0})\to X_{1}\otimes Y_{1}$ is a cofibration. Further, if $f$ or $g$ is a trivial cofibration, so is $f\Box g$.
We say that the symmetric monoidal structure on $\mathcal{M}$ is closed if $\mathcal{M}$ contains an internal hom-object $\mathrm{Map}_{\mathcal{M}}(M,\,N)$ with respect to the monoidal structure $-\otimes-:\mathcal{M}\times\mathcal{M}\to\mathcal{M}$ for any $M,\,N\in\mathcal{M}$. In this case, the push-out product axiom implies that, for any cofibrant object $M$, both those induced functors $-\otimes M:\mathcal{M}\to\mathcal{M}$ and $M\otimes-:\mathcal{M}\to\mathcal{M}$ are left Quillen functors.
Further, we assume that all model categories in this paper are cofibrantly generated. For any class $\mathcal{T}$ of morphisms in a category, ${}^{\boxslash}\mathcal{T}$ (resp. $\mathcal{T}^{\boxslash}$) denotes the class of morphisms having the right (resp. left) lifting property with respect to all morphisms in $\mathcal{T}$. Then ${}^{\boxslash}(\mathcal{T}^{\boxslash})$ is said to be the weakly saturated class of morphisms generated by $\mathcal{T}$. In this paper, we always assume that all model categories are cofibrantly generated: There exist small subsets $\mathcal{I}$ and $\mathcal{J}$ of the set of morphisms of the model category such that the collection of all cofibrations (resp. trivial cofibrations) is the weakly saturated class of morphisms generated by $\mathcal{I}$ (resp. $\mathcal{J}$). In this paper, we call $\mathcal{I}$ (resp. $\mathcal{J}$) the set of generating cofibrations (resp. the set of generating trivial cofibrations). Further, if a cofibrantly generated model category is locally presentable, it is said to be combinatorial. By the theorem of Barwick and Smith [Bar10, Theorem 4.7], for any left proper and combinatorial model category and a small set of homotopy class of morphisms of it, the Bousfield localization exists such that the localized model category is also left proper and combinatorial.
Let $\Delta^{1}$ denote the category with two objects 00 and $1$, and only one non-trivial morphism $0\to 1$. The functor category $\mathrm{Fun}(\Delta^{1},\,\mathcal{M})$ from $\Delta^{1}$ to $\mathcal{M}$ is called the arrow category of $\mathcal{M}$, and $\mathrm{Ar}(\mathcal{M})$ denotes it.
By Hovey [Hov14, Theorem 1.2], the arrow category $\mathrm{Ar}(\mathcal{M})$ has two distinct symmetric monoidal structures derived from $\mathcal{M}$’s: For any two morphisms in $\mathcal{M}$, $f:X_{0}\to X_{1}$ and $g:Y_{0}\to Y_{1}$, one has a commutative square:
${X_{0}\otimes Y_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f\otimes\mathrm{id}}$$\scriptstyle{\mathrm{id}\otimes g}$${X_{1}\otimes Y_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{id}\otimes g}$${X_{0}\otimes Y_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f\otimes\mathrm{id}}$${X_{1}\otimes Y_{1}.}$
One of them is the diagonal (or tensor) product monoidal structure is defined by $f\otimes g$ as the composition
$f\otimes g=(f\otimes\mathrm{id})\circ(\mathrm{id}\otimes g)=(\mathrm{id}\otimes g)\circ(f\otimes\mathrm{id}):X_{0}\otimes Y_{0}\to X_{1}\otimes Y_{1}.$
The other is the push-out product monoidal structure defined by the induced morphism:
$f\Box g:(X_{0}\otimes Y_{1})\amalg_{X_{0}\otimes Y_{0}}(X_{1}\otimes Y_{0})\to X_{1}\otimes Y_{1}.$
[…]
${\mathrm{Ev}_{i}:\mathrm{Ar}(\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\mathcal{M}:L_{i},\,U_{i}\ (i=0,\,1).}$
Namely, those adjoint functors are defined by $L_{0}(X)=(\mathrm{id}:X\to X)$, $U_{0}(X)=(X\to*)$, $L_{1}(X)=(\emptyset\to X)$, and $U_{1}(X)=(\mathrm{id}:X\to X)$. Further, those left adjoints $L_{i}:\mathcal{M}\to\mathrm{Ar}^{\Box}(\mathcal{M})\ (i=0,1)$ are left Quillen and strict monoidal functors, and those right adjoints $U_{i}:\mathcal{M}\to\mathrm{Ar}^{\otimes}(\mathcal{M})\ (i=0,1)$ are right Quillen functors.
Let $\mathrm{Ar}^{\Box}(\mathcal{M})$ (resp. $\mathrm{Ar}^{\otimes}(\mathcal{M})$) denote the symmetric monoidal category with respect to the push-out (resp. tensor) product structure whose monoidal unit is $L_{1}(V)$ (resp. $U_{0}(V)$), where $V$ is the monoidal unit of $\mathcal{M}$. Then $\mathrm{Ar}^{\Box}(\mathcal{M})$ (resp. $\mathrm{Ar}^{\otimes}(\mathcal{M})$) equips a symmetric monoidal model structure with respect to the projective (resp. injective) model structure, which is defined by weak equivalences to be degree-wise weak equivalences and fibrations (resp. cofibrations) to be degree-wise fibrations (resp. degree-wise cofibrations).
Those monoidal model categories $\mathrm{Ar}^{\Box}(\mathcal{M})$ and $\mathrm{Ar}^{\otimes}(\mathcal{M})$ are connected the cokernel-and-kernel adjunction
$\mathrm{cok}:\mathrm{Ar}^{\Box}(\mathcal{M})\rightleftarrows\mathrm{Ar}^{\otimes}(\mathcal{M}):\mathrm{ker},$
which is a Quillen adjunction. The cokernel functor $\mathrm{cok}$ is defined by the sending
$\mathrm{cok}:\mathrm{Ar}^{\Box}(\mathcal{M})$ $\to\mathrm{Ar}^{\otimes}(\mathcal{M})$
$(f:X\to Y)$ $\mapsto(Y\to\mathrm{Coker}(f)),$
and the kernel functor is defined by
$\mathrm{ker}:\mathrm{Ar}^{\otimes}(\mathcal{M})$ $\to\mathrm{Ar}^{\Box}(\mathcal{M})$
$(f:X\to Y)$ $\mapsto(\mathrm{Ker}(f)\to X),$
respectively. In the case that $\mathcal{M}$ is a pointed, left proper, and combinatorial symmetric monoidal model category, the adjunction is lax monoidal Quillen by Hovey’s work (the proof is available at [WY24, Proposition 2.4.3]). Further, if $\mathcal{M}$ is stable, in addition, the adjunction $(\mathrm{cok},\,\mathrm{ker})$ is a Quillen equivalence. (See Hovey [Hov14, Theorem 1.4] or White and Yau [WY24, Proposition 2.3]).
We recall the definition of commutative Smith ideals. A commutative monoid object $A$ of $\mathcal{M}$ is a monoid whose product $\mu_{A}:A\otimes A\to A$ commutes the braiding $\tau_{A,\,A}:A\otimes A\to A\otimes A$, denoting $\mathrm{CAlg}(\mathcal{M})$ the subcategory of commutative objects of $\mathcal{M}$.
Definition 2.1 (c.f. [WY24] Definition 3.3.1).
Let $\mathcal{M}$ be a pointed symmetric monoidal model category. A commutative Smith ideal is a commutative monoid object of the symmetric monoidal model category $\mathrm{Ar}^{\Box}(\mathcal{M})$.
2.2. Homotopically idempotent commutative Smith ideals
We introduce a homotopy analogue of idempotent ideals as follows:
Definition 2.2.
Let $\mathcal{M}$ be a pointed symmetric monoidal model category and $V$ a monoidal unit object. A commutative Smith ideal $j:I\to V$ is homotopically idempotent if it satisfies the following properties:
  • (1)
    The homotopically coCartesian square
[…]
    is also homotopy Cartesian.
  • (2)
    Write $\tilde{j}:I\otimes I\to I\otimes V\to V$ and $\tilde{I}=I\otimes I$. Then $\tilde{I}$ is a cofibrant object and $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ has a homotopy section $\Delta:\tilde{I}\to\tilde{I}\otimes\tilde{I}$ satisfying the associativity:
[…]
  • (3)
    The tensor product $\tilde{I}\otimes\tilde{I}$ is a group like object. That is, the unit morphism $\tilde{I}\otimes\tilde{I}\to\Omega\Sigma(\tilde{I}\otimes\tilde{I})$ is a weak equivalence, where $\Omega$ is the pointed loop functor and $\Sigma$ the suspension functor on the pointed model category $\mathcal{M}$.
Proposition 2.3.
Let $\mathcal{M}$ be a symmetric monoidal model category with a monoidal unit $V$, and $j:\tilde{I}\to V$ a homotopically idempotent commutative Smith ideal of $V$. Then the product $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ is a weak equivalence.
  • proof.
    We show that the homotopy kernel of $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ is contractible. Let $K$ denote the homotopy kernel of $\mu_{\tilde{I}}$. One has a diagram
[…]
    where all squares are homotopy Cartesian, and the upper two squares are homotopy Cartesian. Since weak equivalences are closed under retracts, $\tilde{I}\to\Omega\Sigma\tilde{I}$ and $K\to\Omega\Sigma K$ are weak equivalence. Since $\mu_{\tilde{I}}$ has a homotopy section, the homotopy cokernel $0\amalg_{\tilde{I}\otimes\tilde{I}}\tilde{I}$ is contractible. Therefore, $\Sigma K$ is contractible, implying that $K$ is also contractible.
Remark 2.4.
Since the homotopy cokernel $0\amalg_{\tilde{I}\otimes\tilde{I}}\tilde{I}$ is contractible, the pushout product $\mu_{\tilde{I}}:\tilde{I}\amalg_{\tilde{I}\otimes\tilde{I}}\tilde{I}\to\tilde{I}$ is a weak equivalence, implying that $\mu_{\tilde{j}}:\tilde{j}\Box\tilde{j}\to\tilde{j}$ is a projective equivalence.
Example 2.5.
Let $V$ be a unital commutative ring and $\mathfrak{m}$ an idempotent ideal of $V$. Under the assumption that $\tilde{\mathfrak{m}}=\mathfrak{m}\otimes_{V}\mathfrak{m}$ is a flat $V$-module, the product $\mu_{\mathfrak{m}}:\mathfrak{m}\otimes_{V}\mathfrak{m}\to\mathfrak{m}$ induces isomorphisms $\mu_{\mathfrak{m}}\otimes\mathrm{id}_{\mathfrak{m}}:\tilde{\mathfrak{m}}\otimes_{V}\mathfrak{m}\to\tilde{\mathfrak{m}}$ and $\mathrm{id}_{\mathfrak{m}}\otimes\mu_{\mathfrak{m}}:\mathfrak{m}\otimes_{V}\tilde{\mathfrak{m}}\to\tilde{\mathfrak{m}}$. Hence the composition
$\Delta=(\mu_{\mathfrak{m}}\otimes\mu_{\mathfrak{m}})^{-1}=(\mathrm{id}_{\tilde{\mathfrak{m}}}\otimes\mu_{\mathfrak{m}})^{-1}\circ(\mu_{\mathfrak{m}}\otimes\mathrm{id}_{\mathfrak{m}})^{-1}:\tilde{\mathfrak{m}}\to\tilde{\mathfrak{m}}\otimes_{V}\tilde{\mathfrak{m}}$
is an isomorphism, satisfying the coherence condition in Definition 2.2.
Definition 2.6.
Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal of the monoidal unit $V$. A morphism $f:M\to N$ in $\mathcal{M}$ is an almost weak equivalence if the induced morphism $\tilde{I}\otimes Q(f):\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(N)$ is a weak equivalence, and $\mathcal{M}^{\rm al}$ denotes the Bousfield localization of $\mathcal{M}$ by the class of almost weak equivalences.
Definition 2.7.
An object $M$ of $\mathcal{M}$ is almost local if the induced morphism $f^{*}:\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(N_{2},\,M)\to\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(N_{1},\,M)$ is an isomorphism. Dually, $M$ is almost colocal if, for any homotopically almost weak equivalence $f:N_{1}\to N_{2}$, the induced map $f_{*}:\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(M,\,N_{1})\to\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(M,\,N_{2})$ is an isomorphism.
Definition 2.8.
Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal. An object $E$ is homotopically $I$-closed if the induced morphism
$\mathrm{Map}_{\mathcal{M}}(\tilde{j},\,E):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,E)\to\mathrm{Map}_{\mathcal{M}}(V,\,E)\to E$
is a weak equivalence. An object $E$ is homotopically $I$-firm if $\tilde{j}:\tilde{I}\to V$ induces a weak equivalence
$\tilde{j}\otimes E:\tilde{I}\otimes E\to E.$
Remark 2.9.
In Quillen’s note [Qui96, Definition 5.2], $I$-closed modules said to be closed. However, this paper considers closed monoidal model categories, so we use the terminology $I$-closed objects for given a homotopically idempotent Smith ideal $I$.
We verify the equivalence between almost local objects and homotopically $I$-closed objects.
Proposition 2.10.
Let $\mathcal{M}$ be a pointed symmetric monoidal model category with a monoidal unit $V$ and $j:I\to V$ a homotopically idempotent Smith ideal.
  1. (1)
    For any object $M$, for any object $M$, $\tilde{j}:\tilde{I}\to V$ induces an almost weak equivalence $\tilde{j}:\tilde{I}\otimes M\to V\otimes M\to M$.
  2. (2)
    An object $E$ is almost local if and only if it is homotopically $I$-closed.
  • proof.
    To prove (1), consider $Q(\tilde{j}):\tilde{I}\to Q(V)\to V$ as a factorization, where the first morphism is a cofibration and the second is a trivial fibration. Then, by the unit condition of symmetric monoidal model categories, $\tilde{I}\otimes Q(V)\otimes Q(M)\to\tilde{I}\otimes V\otimes Q(M)$ is a weak equivalence by definition. Additionally, since $\tilde{I}\otimes\tilde{I}\to\tilde{I}\otimes Q(V)$ is a weak equivalence of cofibrant objects, and $-\otimes Q(M)$ is a left Quillen functor, by Ken Brown’s lemma [Hov99, p.6 Lemma 1.1.12], we have that $\tilde{I}\otimes\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(V)\otimes Q(M)$ is also a weak equivalence. Consequently, $\tilde{I}\otimes\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(M)$ is a weak equivalence.
    The if-direction of (2) is clear by the definition of almost weak equivalences. Now, let $E$ be an almost local object. Since the canonical morphism $\tilde{I}\otimes M\to M$ is an almost weak equivalence for any cofibrant object $M\in\mathcal{M}$ and $\tilde{I}$ is cofibrant, we have a chain of isomorphisms: $\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(M,E)\to\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(M,\mathrm{Map}_{\mathcal{M}}(\tilde{I},E))\simeq\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(\tilde{I}\otimes M,E)$. Thus, by the Yoneda lemma, the induced morphism $E\to\mathrm{Map}(\tilde{I},E)$ is a weak equivalence, implying the only-if direction of (2). $\Box$
Theorem 2.11.
Let $\mathcal{M}$ be a pointed symmetric monoidal model category with a monoidal unit object $V$, $j:I\to V$ be a homotopically idempotent commutative Smith ideal and $\mathcal{M}^{\rm al}$ denoting the Bousfield localization by the collection of almost weak equivalences. Then, the internal-hom functor
$\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-):\mathcal{M}\to\mathcal{M}^{\rm al}$
is equivalent to the Bousfield bilocalization functor.
Furthermore, let $\mathcal{M}^{\rm fi}$ denote the homotopically essential image of $\tilde{I}\otimes Q:\mathcal{M}\to\mathcal{M}$. Then, the induced adjunction
$\tilde{I}\otimes Q:\mathcal{M}^{\rm al}\to\mathcal{M}^{\rm fi}:\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$
is a Quillen equivalence. In particular, for any object $E$, the canonical morphism
$E\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,E)$
is an almost weak equivalence.
  • proof.
Let $L^{\rm al}:\mathcal{M}\to\mathcal{M}^{\rm al}$ be another Bousfield localization by almost weak equivalences. By the universality of $L^{\rm al}$, the natural transformation $\mathrm{Id}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ homotopically factors as
[…]
By (2) in Proposition 2.10, the composition $\psi(L^{\rm al})\circ\varphi(L^{\rm al}):L^{\rm al}\to L^{\rm al}\circ L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,L^{\rm al}(-))$ is an equivalence. Therefore, the induced transformation $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,\varphi):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,L^{\rm al}(-))$ is identified with a homotopy section of $\psi:L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$. Further, since the functor $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ is homotopically idempotent, the factorization
$\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,\psi)\circ\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,\varphi):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,L^{\rm al}(-))\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\circ\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$
implies that $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ is a homotopy retraction of $L^{\rm al}$. Hence, the natural transformation $\psi:L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ is a weak equivalence.
We show that the functor $\tilde{I}\otimes Q:\mathcal{M}\to\mathcal{M}$ induced a left Quillen equivalence $\tilde{I}\otimes Q:\mathcal{M}^{\rm al}\to\mathcal{M}^{\rm fi}$. Since the functor $\tilde{I}\otimes Q:\mathcal{M}\to\mathcal{M}$ sends almost weak equivalences to weak equivalences, it factors through $\mathcal{M}^{\rm al}$. By (1) in Proposition 2.10 and the equivalence $\psi:L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$, the unit $\mathrm{Id}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\circ(\tilde{I}\otimes Q)$ induces an equivalence on the homotopy category of $\mathcal{M}^{\rm al}$, implying that it is a left Quillen equivalence. $\Box$
3. Commutative monoid objects in almost mathematics of pointed closed symmetric monoidal model categories
First, we consider the module category over a unital ring. Let $V$ be a commutative unital ring with an idempotent ideal $\mathfrak{m}$. Assume that $\tilde{\mathfrak{m}}$ is a flat $V$-module. In Smith ideal theory language, a $V$-homomorphism $\tilde{j}:\tilde{\mathfrak{m}}\to V$ is regarded a commutative Smith ideal of $V$. For any $V$-algebra $A$, the almost unitalization $A_{!!}$ can be represented by the pushout product
$\tilde{j}\Box(V\to A)=(V\amalg_{\tilde{\mathfrak{m}}}({\tilde{\mathfrak{m}}\otimes A}))\to A,$
which is an almost isomorphism. In Gabber–Ramero’s text [GR03], $A_{!!}$ is defined by to be the source of $\tilde{j}\Box(V\to A)$ without Smith ideal theory. On the view point of Smith ideal theory, we can easily check that the induced ring homomorphism $A_{!!}\to A$ is an almost isomorphism. Indeed, by the associativity of pushout products,
$(\tilde{\mathfrak{m}}\otimes_{V}{A}_{!!}\to\tilde{\mathfrak{m}}\otimes_{V}A)\simeq L_{1}(\tilde{\mathfrak{m}})\Box(\tilde{j}\Box(V\to A))\to(L_{1}(\tilde{\mathfrak{m}})\Box\tilde{j})\Box(V\to A)\simeq\mathrm{id}_{\tilde{\mathfrak{m}}}\Box(V\to A)\simeq\mathrm{id}_{\tilde{\mathfrak{m}}\otimes_{V}A}$
is an isomorphism between arrows, implying that those sources $\tilde{\mathfrak{m}}\otimes_{V}A_{!!}$ and $\tilde{\mathfrak{m}}\otimes_{V}A$ are isomorphic. We generalize those properties to model categories.
3.1. Almost unitalization functors of commutative algebra objects of symmetric monoidal model categories
In general the adjunction $\mathrm{Sym}:\mathcal{M}\rightleftarrows\mathrm{CAlg}(\mathcal{M}):U$ does not induces a model structure on $\mathrm{CAlg}(\mathcal{M})$ but it has the semi-model structure defined in [Spi01, Definition 2.3] or [Whi22, Definition 3.9]. Semi-model structure only needs that trivial cofibrations with cofibrant domains have the left lifting property with respect to all fibrations, and only maps with cofibrant domains factor into a trivial cofibration followed by a fibration. If all objects of a semi-model category are cofibrant, it is just a model category (See [Whi22, Definition 3.9]). Left and right Quillen functors between semi-model categories are defined as the same as ones of model categories.
For any class of morphisms $\mathcal{I}$, let $\mathrm{Cell}(\mathcal{I})$ denote the class of morphisms that is a transfinite composition of pushouts of the element of $\mathcal{I}$, and an object of $\mathrm{Cell}(\mathcal{I})$ is said to be a relative $\mathcal{I}$-cell complex.
Let $\mathcal{M}$ is cofibrantly generated with a set of generating cofibrations $\mathcal{I}$ and a set of generating trivial cofibrations $\mathcal{J}$. By White [Whi22, Corollary 3.8], if the domains of $\mathcal{I}$ (resp. $\mathcal{J}$) is small relative to $\mathrm{Cell}(\mathcal{I}\otimes\mathcal{M})$ (resp. $\mathrm{Cell}(\mathcal{J}\otimes\mathcal{M})$), then the category of commutative $R$-algebras has a semi-model structure inherited by $\mathcal{M}$ for any commutative monoid object $R$.
We recall the definition of lax monoidal Quillen adjunctions:
Definition 3.1 ([SS03] Definition 3.6).
Let $F:\mathcal{M}\rightleftarrows\mathcal{N}:U$ be a Quillen adjunction of monoidal model categories. Then we say that $(F,\,U)$ is lax Quillen monoidal if it satisfies the following properties:
  • (1)
    The right Quillen functor $U$ is lax monoidal.
  • (2)
    For any cofibrant objects $M$ and $N$ in $\mathcal{M}$, the induced morphism
    $\nabla_{M,\,N}:F(M\otimes N)\to F(M)\otimes F(N)$
    by the canonical morphism $\Delta_{M,\,N}:M\otimes N\to U(F(M))\otimes U(F(N))\to U(L(M)\otimes L(N))$ is a weak equivalence.
  • (3)
    For any cofibrant replacement $q(\mathbf{1}_{\mathcal{M}}):Q(\mathbf{1}_{\mathcal{M}})\to\mathbf{1}_{\mathcal{M}}$ of the monoidal units $\mathbf{1}_{\mathcal{M}}$ of $\mathcal{M}$ and $\mathbf{1}_{\mathcal{N}}$ of $\mathcal{N}$, the composition
    $F(Q(\mathbf{1}_{\mathcal{M}}))\to F(\mathbf{1}_{\mathcal{M}})\to\mathbf{1}_{\mathcal{N}}$
    is a weak equivalence.
Proposition 3.2.
Let $j:I\to M$ be a homotopically idempotent commutative Smith ideal. Then the Quillen equivalence
$\tilde{I}\otimes Q:\mathcal{M}^{\rm al}\rightleftarrows\mathcal{M}^{\rm fi}:L^{\rm al}$
is lax Quillen monoidal.
  • proof.
    Since the class almost weak equivalences between cofibrant objects is closed under tensor product, the localization $L^{\rm al}:\mathcal{M}\to\mathcal{M}^{\rm al}$ is lax monoidal. Furthermore, given any pair of cofibrant objects $M$ and $N$, the product $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ induces a weak equivalence
    $(\tilde{I}\otimes Q(M))\otimes(\tilde{I}\otimes Q(N))\simeq(\tilde{I}\otimes M)\otimes(\tilde{I}\otimes N)\to\tilde{I}\otimes(M\otimes N).$
    Furthermore, since $Q(M)\otimes Q(N)$ is cofibrant, $Q(M)\otimes Q(N)\to M\otimes N$ factors through the cofibrant replacement $Q(M\otimes N)$, implying that Condition (2) of Definition 3.1 holds. Condition (3) is obvious. $\Box$
Corollary 3.3.
Let $j:I\to M$ be a homotopically idempotent commutative Smith ideal. The mapping functor $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-):\mathcal{M}\to\mathcal{M}^{\rm al}$ is lax monoidal, and the Quillen equivalence $(\tilde{I}\otimes Q,\,\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,R(-)))$ induces on their commutative monoid objects
$\tilde{I}\otimes Q:\mathrm{CAlg}(\mathcal{M}^{\rm al})\rightleftarrows\mathrm{CAlg}(\mathcal{M}^{\rm fi}):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$
between semi-model categories. $\Box$
Now, for any commutative monoid object $A$ of $\mathcal{M}$, we define a monoid object $A_{!!}$ and the induced morphism $A_{!!}\to A$ is an almost weak equivalence under some mild conditions.
Proposition 3.4.
Let $\mathcal{M}$ be a pointed closed symmetric monoidal model category with a monoidal unit object $V$ and $A$ a commutative algebra object of $\mathcal{M}$. Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal of $V$. For any monoid morphism $f:A\to B$ of commutative monoid objects, the induced morphism
$\tilde{j}\Box f:(V\otimes A)\amalg_{\tilde{I}\otimes A}(\tilde{I}\otimes B)\to B$
is a monoid morphism.
  • proof.
    Replacing $V$ and $A$ and considering $\mathrm{CAlg}(\mathcal{M})_{A/}$, it is sufficient to prove that
    $\tilde{j}\Box L_{1}(A):A_{!!}=V\amalg_{\tilde{I}}(\tilde{I}\otimes A)\to A$
    is a monoid morphism between commutative monoid objects. The monoid structure on $A_{!!}$ is induced by those morphisms
    $\eta_{A}:V\otimes A\to A\ \text{ and }\ \tilde{j}\otimes\mathrm{Id_{A}}:\tilde{I}\otimes A\to V\otimes A\to A,$
    and the induced morphism
    $\mu_{\tilde{I}}\otimes\mu_{A}:(\tilde{I}\otimes A)\otimes(\tilde{I}\otimes A)\to(\tilde{I}\otimes A).$
    Then those morphisms induces $\eta_{A_{!!}}:V\otimes A_{!!}\to A_{!!}$ and $\mu_{!!}:A_{!!}\otimes A_{!!}\to A_{!!}$ satisfy the associativity, the unitality, and the commutativity inherited by all of $V$, $\tilde{I}$, and $\tilde{I}\otimes{A}$. Hence, $A_{!!}$ has a commutative monoid structure, and the square $(\tilde{I}\to V)\to(\tilde{I}\otimes A\to A)$ is a morphism of commutative Smith ideals, implying that $A_{!!}\to A$ is a morphism of commutative monoid objects. $\Box$
Proposition 3.5.
Let $\mathcal{M}$ be a pointed closed symmetric monoidal model category with a monoidal unit object $V$ and $A$ a commutative algebra object of $\mathcal{M}$. Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal of $V$. If $0\to V\to A$ or $V\to A$ is a cofibration, then the pushout product
$(\tilde{j}\Box L_{1}(A):A_{!!}\to A)\to(\mathrm{Id}_{A})$
is an almost weak equivalence, where we write
$A_{!!}=\mathrm{Ev}_{0}((\tilde{j}:\tilde{I}\to V)\Box(L_{1}(A):V\to A))=V\amalg_{\tilde{I}}(\tilde{I}\otimes A)).$
  • proof.
    Assume that $0\to A$ is a cofibration. By definition of homotopy pushout, let $\tilde{I}\to(\tilde{I}\otimes A)^{\prime}\to\tilde{I}\otimes A$ be a factorization of a cofibration to a trivial fibration, then the induced morphisms
    $Q(V)\amalg_{\tilde{I}}(\tilde{I}\otimes A)^{\prime}\to Q(A_{!!})\to A_{!!}$
    are weak equivalences. Therefore $\tilde{I}\otimes(\tilde{I}\otimes A)^{\prime}\to\tilde{I}\otimes Q(A_{!!})$ is a weak equivalence. Since $(\tilde{I}\otimes A)^{\prime}\to\tilde{I}\otimes A$ is a weak equivalence between cofibrant objects, $\tilde{I}\otimes(\tilde{I}\otimes A)^{\prime}\to\tilde{I}\otimes(\tilde{I}\otimes A)\to\tilde{I}\otimes A$ is a composition of weak equivalences. Therefore $\tilde{I}\otimes Q(A_{!!})\to\tilde{I}\otimes A$ is a weak equivalence.
    Next, we assume that $V\to A$ is a cofibration. Then $\tilde{I}\otimes A$ is cofibrant. Therefore $\tilde{I}\otimes A\to A$ factors as
    $\tilde{I}\otimes A\to Q(A_{!!})\to Q(A)\to A,$
    inducing a tower of morphisms
    $\cdots\to\tilde{I}^{\otimes 3}\otimes Q(A)\to\tilde{I}^{\otimes 3}\otimes A\to\tilde{I}^{\otimes 2}\otimes Q(A_{!!})\to\tilde{I}^{\otimes 2}\otimes Q(A)\to\tilde{I}\otimes A\to\tilde{I}\otimes Q(A_{!!})\to\tilde{I}\otimes Q(A).$
    Since $\tilde{I}^{\otimes 2}\otimes Q(A_{!!})\to\tilde{I}^{\otimes 2}\otimes Q(A)\to\tilde{I}^{\otimes 2}\otimes A\to\tilde{I}\otimes Q(A_{!!})$ and $\tilde{I}^{\otimes 2}\otimes Q(A)\to\tilde{I}\otimes A\to\tilde{I}\otimes Q(A_{!!})\to\tilde{I}\otimes Q(A)$ are weak equivalences, we obtain that $\tilde{I}\otimes Q(A_{!!})\to\tilde{I}\otimes Q(A)$ is also a weak equivalence. $\Box$
We verify that, for any commutative monoid object $A$, the internal-hom object $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)$ has a commutative monoid structure on $\mathcal{M}$ inherited by one of $A$ and the homotopy section $\Delta:\tilde{I}\to\tilde{I}\otimes\tilde{I}$ satisfying the coherent conditions in Definition 2.2. Let $\mathrm{ev}_{\tilde{I}}:\tilde{I}\otimes\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)\to A$ denote the evaluation morphism induced by the identity on $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)$. Then the composition
$\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)\otimes\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)\otimes(\tilde{I}\otimes\tilde{I})\cong(\tilde{I}\otimes\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A))\otimes(\tilde{I}\otimes\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A))\overset{\mathrm{ev}_{\tilde{I}}\otimes\mathrm{ev}_{\tilde{I}}}{\to}A\otimes A\overset{\mu_{A}}{\to}A$
determines a canonical monoid structure
$\mu_{\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)}:\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)\otimes\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)\to\mathrm{Map}_{\mathcal{M}}(\tilde{I}\otimes\tilde{I},\,A)\overset{\Delta^{*}}{\to}\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A),$
satisfying the associativity and unitality due to Condition (3) in Definition 2.2.
In the case that the monoidal structure on $\mathcal{M}$ is closed, both the symmetric monoidal structures on the allow category $\mathcal{M}$ are closed by Hovey [Hov14, Theorem 1.2]. For any couple of morphisms $(f:X_{0}\to X_{1},\,g:Y_{0}\to Y_{1})$ in $\mathcal{M}$, the internal-hom objects are determined as follows: The internal-hom object $\mathrm{Map}_{\mathcal{M}}^{\otimes}(f,\,g)$ of the diagonal product monoidal structure is defined by
$\mathrm{Map}^{\otimes}(f,\,g):\mathrm{Map}_{\mathcal{M}}(X_{0},\,Y_{0})\times_{\mathrm{Map}_{\mathcal{M}}(X_{0},\,Y_{1})}\mathrm{Map}_{\mathcal{M}}(X_{1},\,Y_{1})\to\mathrm{Map}_{\mathcal{M}}(X_{1},\,Y_{1})$
and $\mathrm{Map}^{\Box}(f,\,g)$ of the pushout monoidal structure is defined by
$\mathrm{Map}^{\Box}(f,\,g):\mathrm{Map}_{\mathcal{M}}(X_{1},\,Y_{0})\to\mathrm{Map}_{\mathcal{M}}(X_{0},\,Y_{0})\times_{\mathrm{Map}_{\mathcal{M}}(X_{0},\,Y_{1})}\mathrm{Map}_{\mathcal{M}}(X_{1},\,Y_{1}).$
This section’s main result is the following.
Theorem 3.6.
Let $\mathcal{M}$ be a pointed closed symmetric monoidal model category with a monoidal unit object $V$ and homotopically idempotent commutative Smith ideal $j:I\to V$. Assume that $\mathcal{M}$ is left proper and combinatorial. Then the Quillen adjunction
$\tilde{j}\Box(-):\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}))\rightleftarrows\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M})):\mathrm{Map}^{\Box}(\tilde{j},\,-)$
induces a Quillen equivalence
$\tilde{j}\Box(-):\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al}))\rightleftarrows\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al})):\mathrm{Map}^{\Box}(\tilde{j},\,-),$
which is equivalent to the trivial Quillen equivalence
$L_{0}\circ\mathrm{Ev}_{1}:\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al}))\rightleftarrows\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al})):U_{1}\circ\mathrm{Ev}_{0}.$
  • proof.
    Since the subcategory $\mathrm{CAlg}(\mathcal{M})$ is closed under homotopy limits, for any morphism $f:A\to B$ of commutative monoid objects, $\mathrm{Map}^{\Box}(\tilde{j},\,f):A\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)\times_{\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,B)}B$ is a morphism of monoid objects by Corollary 3.3. The functor $\mathrm{Map}^{\Box}(\tilde{j},-):\mathrm{Ar}(\mathcal{M})\to\mathrm{Ar}(\mathcal{M})$ can be restricted on $\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}))$.
By Proposition 3.5, $j\Box\mathrm{Id}\to\mathrm{id}_{Ev_{1}(-)}=L_{0}(\mathrm{Ev}_{1}((-)))$ is a projective equivalence in $\mathrm{CAlg}(\mathcal{M}^{\rm al})$. Therefore, their Quillen right adjoints $\mathrm{Map}^{\Box}(\tilde{j},\,-)$ and $U_{1}\circ\mathrm{Ev}_{0}$ are weakly equivalent on $\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al}))$. $\Box$
We obtain an analogue of Gabber–Ramero [GR03, p.22, Proposition 2.2.29] as avv corollary of Theorem 3.6.
Corollary 3.7.
Let $\mathcal{M}$ be a left proper and combinatorial pointed closed symmetric monoidal model category with a monoidal unit object $V$ and a homotopically idempotent commutative Smith ideal $j:I\to V$. Then the functor $(-)_{!!}=\mathrm{Ev}_{0}(\tilde{j}\Box L_{1}(-)):\mathrm{CAlg}(\mathcal{M})\to\mathrm{CAlg}(\mathcal{M})$ is a left Quillen functor, whose right adjoint is the composition
$\mathrm{Ev}_{1}(\mathrm{Map}^{\Box}(\tilde{j},\,U_{0}(-))):\mathrm{CAlg}(\mathcal{M})\to\mathrm{CAlg}(\mathcal{M})$
Furthermore, the left Quillen functor induces a left Quillen equivalence
$(-)_{!!}:\mathrm{CAlg}(\mathcal{M}^{\rm al})\to\mathrm{CAlg}(\mathcal{M})_{!!},$
where $\mathrm{CAlg}(\mathcal{M})_{!!}$ denotes the homotopically essential image of the endofunctor
$\mathrm{Ev}_{0}(\tilde{j}\Box L_{1}(-)):\mathrm{CAlg}(\mathcal{M})\to\mathrm{CAlg}(\mathcal{M}).$
  • proof.
    Since all of $L_{1}$, $\tilde{j}\Box(-)$, and $\mathrm{Ev}_{0}$ are left Quillen functors, the composition $(-)_{!!}$ is also a left Quillen functor, and its right adjoint is the composition of their right adjoint functors
    $\mathrm{Ev}_{1}\circ\mathrm{Map}^{\Box}(\tilde{j},\,-)\circ U_{0}:\mathrm{CAlg}(\mathcal{M})\to\mathrm{CAlg}(\mathcal{M}).$
    By Theorem 3.6, one has a weak equivalence $L^{\rm al}\circ(-)_{!!}\circ L^{\rm al}\simeq L^{\rm al}$, and by Proposition 3.5, the endofunctor $(-)_{!!}:\mathrm{CAlg}(\mathcal{M})\to\mathrm{CAlg}(\mathcal{M})$ is weakly idempotent. Hence, those compositions
    $L^{\rm al}\circ(-)_{!!}:\mathrm{CAlg}(\mathcal{M}^{\rm al})\to\mathrm{CAlg}(\mathcal{M})_{!!}\to\mathrm{CAlg}(\mathcal{M}^{\rm al})$
    and
    $(-)_{!!}\circ L^{\rm al}:\mathrm{CAlg}(\mathcal{M})_{!!}\to\mathrm{CAlg}(\mathcal{M}^{\rm al})\to\mathrm{CAlg}(\mathcal{M})_{!!}$
    are equivalent to the identity functors, respectively. $\Box$
4. Bousfield bilocalization of stable model categories
A pointed model category is stable if its homotopy category is a triangulated category. In a stable model category, a homotopy Cartesian square is also homotopy coCartesian and vice versa. Further, in this section, we consider the case that stable monoidal model categories are generated by the monoidal unit objects. Those such categories said to be monogenic, and we explain in Definition 4.1. In this section, we prove that Bousfield bilocalization functors determine homotopically idempotent Smith ideals under the condition that those model categories are stable closed symmetric monoidal and monogenic. Furthermore, we show that homotopically closed objects with respect to those determined Smith ideals by Bousfield bilocalization are homotopically coincide with local objects.
4.1. Monogenic model categories
Following MacLane [Mac88, Chapter V, Section 7, p.127], we recall the definition of generators of categories: Let $\mathcal{M}$ be a category and $\mathcal{G}$ a class of small objects of $\mathcal{M}$. The set $\mathcal{G}$ is a generator of $\mathcal{M}$ if for any pair of morphisms $h,\,h^{\prime}:X\to Y$ in $\mathcal{M}$, $h\neq h^{\prime}$ implies that there exists $S\in\mathcal{G}$ and a morphism $f:S\to X$ such that $h\circ f\neq h^{\prime}\circ f$. As Hovey’s book [Hov99, p.183, Definition 7.2.1], we define weakly generators of model categories.
Definition 4.1 (c.f. [HPS97] p.7).
Let $\mathcal{M}$ be a model category with a class $\mathcal{G}$ of objects of $\mathcal{M}$. We say that $\mathcal{G}$ weakly generates $\mathcal{M}$ if the equivalence class of $\mathcal{G}$ generates the homotopy category of $\mathcal{M}$. A symmetric monoidal model category weakly generated by the small monoidal unit is said to be monogenic.
By the following proposition, for stable model categories, the above definition coincides with [Hov99, p.183, Definition 7.2.1].
Proposition 4.2.
Let $\mathcal{M}$ be a pointed closed symmetric monoidal model category with a class $\mathcal{G}$ of objects of $\mathcal{M}$. Assume that $\mathcal{G}$ generates $\mathcal{M}$. Then an object $M$ is contractible if and only if the internal hom-object $\mathrm{Map}_{\mathcal{M}}(S,\,M)$ is contractible for any $S\in\mathcal{G}$.
  • proof.
    The if-direction is clear. Let $M$ be an object of $\mathcal{M}$ such that the internal-hom object $\mathrm{Map}_{\mathcal{M}}(S,\,M)$ is contractible for any $S\in\mathcal{G}$. Since $\mathcal{G}$ generates $\mathcal{M}$, for any object $N$ and morphism $f:N\to M$, the condition that $f$ is not null-homotopic implies that there exists an object $S\in\mathcal{G}$ and a morphism $i:S\to N$ and $f\circ i:S\to M$ is not null-homotopic, contradicting the generating condition. $\Box$
4.2. Symmetric monoidal bilocalization on stable model categories
Definition 4.3.
Let $\mathcal{M}$ be a model category. A functor $F:\mathcal{M}\to\mathcal{N}$ is said to be a Bousfield bilocalization if it is a Bousfield localization admitting fully faithful left adjoint and fully faithful right adjoint such that all of those functors are Quillen functors.
We consider a Bousfield bilocalization $F:\mathcal{M}\to L(\mathcal{M})$ and let denote $F_{*}$ its fully faithful left adjoint and $U$ its fully faithfully right
${F:\mathcal{M}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces L(\mathcal{M}).}$$\scriptstyle{U}$$\scriptstyle{F_{*}}$
Further, we assume that the Quillen adjunction
$F_{*}:L(\mathcal{M})\rightleftarrows\mathcal{M}:F$
is lax Quillen monoidal. In particular, for any $M$ and $N$ of $\mathcal{M}$, the induced morphism
$\nabla:F_{*}(M)\otimes F_{*}(N)\to F_{*}(M\otimes N)$
is a weak equivalence. Then the counit $c:F_{*}\circ F\to\mathrm{Id}_{\mathcal{M}}$ induces a commutative Smith ideal $c_{V}:F_{*}(F(Q(V)))\to V$. Let $j:I\to V$ denote a cofibrant replacement of the homotopy image of $c_{V}$. We will prove that $I$ is homotopically idempotent.
Lemma 4.4.
Let $F:\mathcal{M}\to L(\mathcal{M})$ be a Bousfield bilocalization of stable symmetric monoidal model category, and $F_{*}$ denote its fully faithful left adjoint. Assume that the Quillen adjunction
$F_{*}:L(\mathcal{M})\rightleftarrows\mathcal{M}:F$
is lax Quillen monoidal. Let $j:I\to V$ be a homotopy image object of $q(V)\circ c_{Q(V)}:F_{*}(F(Q(V))\to Q(V)\to V$. Then all of the canonical morphisms in the diagram
$I\otimes I\rightarrow F_{*}(F(Q(V))\otimes I\rightarrow F_{*}(F(Q(V))\otimes F_{*}(F(Q(V))\rightarrow F_{*}(F(Q(V)))\leftarrow I$
are weak equivalences.
  • proof.
    Let $q(V):Q(V)\to V$ be a cofibrant replacement of $V$. Since the Bousfield localization $F$ is a left and right Quillen functor, $F(Q(V))\to F(V)$ is a trivial fibration, and $F(Q(V))$ is a cofibrant object of $L(\mathcal{M})$. Furthermore, since $F_{*}$ is lax Quillen monoidal, $F(Q(V))$ is equivalent to a monoidal unit of $L(\mathcal{M})$. Therefore, all morphisms in
    $F(Q(V))\otimes F(Q(V))\to F(V)\otimes F(Q(V))\to F(Q(V))\to F(V)$
    are weak equivalences by Ken Brown’s lemma [Hov99, p.6 Lemma 1.1.12]. In particular,
    $\mu:F(Q(V))\otimes F(Q(V))\to F(Q(V))$
    is a weak equivalence between cofibrant objects. Since $F_{*}$ is a left Quillen functor, by Ken Brown’s lemma [Hov99, p.6 Lemma 1.1.12] again,
    $F_{*}(\mu):F_{*}(F(Q(V))\otimes F(Q(V)))\to F_{*}(F(Q(V)))$
    is a weak equivalence. Note that the product
    $\mu_{F_{*}(F(Q(V)))}:F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))\to F_{*}(F(Q(V)))$
    is homotopic to the pushout of $\nabla:F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V))\to F_{*}(F(Q(V))\otimes F(Q(V)))$ along the morphism $F_{*}(F(\mu)):F_{*}(F(Q(V))\otimes F(Q(V)))\to F_{*}(F(Q(V)))$. By the left properness of $\mathcal{M}$ implies that $\mu:F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))\to F_{*}(F(Q(V)))$ is a weak equivalence.
By the stability of $\mathcal{M}$, the canonical morphism $(j:I\to V)\to(q(V)\circ c_{Q(V)}:F_{*}(F(Q(V)))\to V)$ is a weak equivalence in the projective model structure. Since $I\to F_{*}(F(Q(V)))$ is a weak equivalence of cofibrant objects, by Ken Brown’s lemma again, both morphisms in $I\otimes I\rightarrow F_{*}(F(Q(V))\otimes I\rightarrow F_{*}(F(Q(V))\otimes F_{*}(F(Q(V))$ are weak equivalences. $\Box$
Proposition 4.5.
Let $F:\mathcal{M}\to L(\mathcal{M})$ be a Bousfield bilocalization of stable symmetric monoidal model category, and $F_{*}$ denote its fully faithful left adjoint. Assume that the Quillen adjunction
$F_{*}:L(\mathcal{M})\rightleftarrows\mathcal{M}:F$
is lax Quillen monoidal. Then the counit $c:F_{*}\circ F\to\mathrm{Id}_{\mathcal{M}}$ induces determines a homotopically idempotent commutative Smith ideal $j:I\to V$ of a monoidal unit object $V$, where $I$ is a cofibrant replacement of the homotopy image of $q(V)\circ c_{Q(V)}:F_{*}(F(Q(V))\to Q(V)\to V$.
  • proof.
    Since $\mathcal{M}$ is stable, the unit morphism $c_{V}\to j$ is a projective weak equivalence in the arrow category of $\mathrm{Ar}(\mathcal{M})$. Condition (1) of Definition 2.2 is already satisfied. By Lemma 4.4, condition (2) in Definition 2.2 holds. By the stability of $\mathcal{M}$, the right exact functor $I\otimes(-)$ is also left exact, implying that the induced morphism $\tilde{I}\otimes I\to\tilde{I}$ is a weak equivalence. $\Box$
Proposition 4.6.
Let $\mathcal{M}$ be a pointed closed symmetric monoidal model with a monoidal unit $V$ and $F:\mathcal{M}\to\mathcal{M}$ a symmetric monoidal Bousfield bilocalization, and $F_{*}:\mathcal{M}\to\mathcal{M}$ denote the homotopically fully faithful left adjoint. Assume that the Quillen adjunction
$F_{*}:L(\mathcal{M})\rightleftarrows\mathcal{M}:F$
is lax Quillen monoidal. Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal determined by the counit $c:F_{*}\circ F\to\mathrm{Id}$. Then
$F(I\otimes Q(M))\to F(M)$
is a weak equivalence for any object $M$.
  • proof.
    Let $j:I\to F_{*}(F(V))$ be a cofibrant replacement of the homotopy image of the counit $c_{V}:F_{*}(F(V))\to V$. Then $F(j):F(I)\to\mathrm{Im}(c_{V})$ is a trivial fibration. Therefore, $F(I)\to F(F_{*}(F(V))$ is a weak equivalence. Since $F_{*}$ is fully faithful, the natural transformation $F(c):F\circ F_{*}\circ F\to F$ is a weak equivalence. Therefore, $F(I)\to F(F(V))$ is also a weak equivalence. Therefore
    $F(\tilde{I})\to F(I)\otimes F(I)\to F(F_{*}(F(Q(V))))\otimes F(F_{*}(F(Q(V))))\\ \to F(Q(V))\otimes F(Q(V))\to F(V)\otimes F(Q(V))\to F(V)\otimes F(V)\to F(V)$
    is a weak equivalence and $F(\tilde{I}\otimes Q(M))\to F(\tilde{I})\otimes F(Q(M))\to F(V)\otimes F(Q(M))\to F(Q(M))$ is a weak equivalence for any object $M\in\mathcal{M}$. Since $F$ is a right Quillen functor, for any cofibrant replacement $Q(M)\to M$, $F(Q(M))\to F(M)$ is a trivial fibration, giving us the conclusion. $\Box$
Theorem 4.7.
Let $\mathcal{M}$ be a stable closed symmetric monoidal model category and $F:\mathcal{M}\to L(\mathcal{M})$ a Bousfield bilocalization, and $F_{*}:\mathcal{M}\to\mathcal{M}$ denote the homotopically fully faithful left adjoint. Assume that $\mathcal{M}$ is monogenic and the Quillen adjunction
$F_{*}:L(\mathcal{M})\rightleftarrows\mathcal{M}:F$
is lax Quillen monoidal. Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal determined as a result of Proposition 4.5. Let $U$ denote a right Quillen adjoint of $F$. Then all of the morphisms of the diagram
[…]
are weak equivalences for any fibrant object $M$ of $\mathcal{M}$.
  • proof.
    Let $j:\tilde{I}\to V$. For any fibrant object $M$ of $\mathcal{M}$, one has a commutative diagram
[…]
    where all vertical maps are isomorphisms. Since $M$ is fibrant, all of the lower arrows induced by weak equivalences in Lemma 4.4 are isomorphisms. Therefore both $\varphi_{*}$ and $\psi_{*}$ are isomorphisms. Since $V$ generates $\mathcal{M}$, both of the homotopy cokernels of $\varphi$ and $\psi$ are contractible, implying that they are weak equivalences. $\Box$
5. Operadic non-unital algebra objects of stable symmetric monoidal model categories
5.1. Operadic Smith ideals
Remark 5.1.
The definitions of colors, symmetric sequences, and colored operads in symmetric monoidal categories can be found in Yau’s book, specifically in [Yau16, Chapter 9 and 11]. Colored operads and maps colored operads are defined in [Yau16, p.176, Definition 11.2.1] and [Yau16, p.180, Definition 11.2.12], respectively. Additionally, the material of algebras over colored operads is explained in [Yau16, p.217, Definition 13.2.1].
Further, in the case that the monoidal structure on $\mathcal{M}$ is closed, algebras over $\mathcal{O}$ are determined by operad maps from $\mathcal{O}$ to endomorphism operads [Yau16, p.226, Definition 13.8.1]. For any $\mathfrak{C}$-colored object $A=\{A_{c}\}_{c\in\mathfrak{C}}$ of $\mathcal{M}$, the endomorphism operad $\mathbf{End}_{\mathcal{M}}(A)$ is defined by
$\mathbf{End}_{\mathcal{M}}(A)(c_{1},\ldots,c_{n};d)=\mathrm{Map}_{\mathcal{M}}(A_{c_{1}}\otimes\cdots\otimes A_{c_{n}},A_{d}).$
Giving a map $\theta:\mathcal{O}\to\mathbf{End}_{\mathcal{M}}(A)$ of $\mathfrak{C}$-colored operads determines an $\mathcal{O}$-algebra structure on $A$. (See [Yau16, p.231, Proposition 13.9.1]).
Definition 5.2.
Let $\mathcal{M}$ be a symmetric monoidal closed (model) category and $\mathcal{O}$ a colored operad in $\mathcal{M}$. Then we write $\overrightarrow{\mathcal{O}^{\Box}}=L_{1}(\mathcal{O})$ and $\overrightarrow{\mathcal{O}^{\otimes}}=L_{0}(\mathcal{O})$. Since $L_{1}$ (resp. $L_{0}$) is strict monoidal (See [WY24, Definition 3.1.1 (7)]), $\overrightarrow{\mathcal{O}^{\Box}}$ (resp. $\overrightarrow{\mathcal{O}^{\otimes}}$) is a colored operad in $\mathrm{Ar}^{\Box}(\mathcal{M})$ (resp. $\mathrm{Ar}^{\otimes}(\mathcal{M})$). Let $\mathrm{Alg}_{\overrightarrow{\mathcal{O}^{\Box}}}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ (resp. $\mathrm{Alg}_{\overrightarrow{\mathcal{O}^{\otimes}}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))$) denote the category of $\mathcal{O}^{\Box}$-algebra objects of $\mathrm{Ar}^{\Box}(\mathcal{M})$ (resp. $\mathcal{O}^{\otimes}$-algebra objects of $\mathrm{Ar}^{\otimes}(\mathcal{M})$).
Operadic Smith ideals defined as the following:
Definition 5.3 ([WY24] Definition 3.3.1).
Let $\mathcal{M}$ be a symmetric monoidal closed (model) category and $\mathcal{O}$ a colored operad in $\mathcal{M}$. Then an $\mathcal{O}$-Smith ideal is an object of $\mathrm{Alg}_{\overrightarrow{\mathcal{O}^{\Box}}}(\mathrm{Ar}^{\Box}(\mathcal{M}))$.
5.2. Non-unital algebras from the (operadic) Smith ideal theory viewpoint
We recall a brief review of non-unital rings, which are rings without assuming the existence of the multiplicative unit $1$. Any unital ring is regarded as a non-unital ring by forgetting the existence of $1$. An augmentation $\varepsilon_{A}:A\to V$ is a ring homomorphism that the unit homomorphism $V\to A$ is a section of $\varepsilon_{A}$, and a $V$-algebra with an augmentation is called an augmented $V$-algebra. Let $\mathrm{Alg}_{V//V}$ denotes the category of augmented $V$-algebras and $\mathrm{Alg}^{\rm nu}_{V}$ the category of non-unital $V$-algebras. It is well-known that there exists a categorical equivalence between $\mathrm{Alg}^{\rm nu}_{V}$ and $\mathrm{Alg}_{V//V}$ as follows: For any $V$-algebra $A$, the direct sum $V\oplus A$ has a canonical unital ring structure defined by
$(m,\,a)\cdot(n,\,b)=(mn,\,na+mb+ab)$
for $m,\,n\in V$ and $a,\,b\in A$. It is easily checked that the functor the adjunction
$V\oplus(-):\mathrm{Alg}^{\rm nu}_{V}\rightleftarrows\mathrm{Alg}_{V//V}:\mathrm{Ker}(\epsilon_{(-)}:-\to V)$
is a pair of categorical equivalences, where $\mathrm{Ker}(\varepsilon_{(-)}:-\to V)$ is the kernel functor of augmentations.
From the Smith ideal theory viewpoint, the above categorical equivalence is a special case of cokernel-and-kernel correspondence between the arrow categories. We consider the case abelian categories. Let $\mathcal{M}$ be a symmetric closed monoidal abelian category and $\mathrm{Ar}^{\rm im}(\mathcal{M})$ denote the full subcategory spanned those objects $f:R\to S$ such that the unit $f\to\mathrm{ker}(\mathrm{cok}(f))$ is an isomorphism. Then the cokernel functor $\mathrm{Ar}^{\rm im}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$ is a categorical equivalence.
Proposition 5.4.
Let $\mathcal{M}$ be a locally presentable abelian symmetric monoidal category. Then, the arrow category $\mathrm{Ar}(\mathcal{M})$ is also locally presentable and there exist reflective localization functors $L_{\rm im}:\mathrm{Ar}^{\Box}(\mathcal{M})\to\mathrm{Ar}^{\rm im}(\mathcal{M})$ by the unit $\mathrm{id}\to\ker\circ\mathrm{coker}$ and $L_{\rm coim}:\mathrm{Ar}^{\otimes}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$ by the counit $\mathrm{coker}\circ\ker\to\mathrm{id}$ such that the adjunction
$\mathrm{cok}:\mathrm{Ar}^{\Box}(\mathcal{M})\rightleftarrows\mathrm{Ar}^{\otimes}(\mathcal{M}):\mathrm{ker}$
induces categorical equivalences
$\mathrm{cok}:\mathrm{Ar}^{\rm im}(\mathcal{M})\rightleftarrows\mathrm{Ar}^{\rm coim}(\mathcal{M}):\mathrm{ker}.$
  • proof.
    Since any abelian category is binormal, $f\to\mathrm{im}(f)$ is an isomorphism if and only if $f$ is a monomorphism. By Adámek–Rosický [AR94, p.44, Corollary 1.5.4], the arrow category is locally presentable, and it admits reflective localization. By the definitions of those functors: $\mathrm{im}=\mathrm{ker}\circ\mathrm{cok}$ and $\mathrm{coim}=\mathrm{cok}\circ\mathrm{ker}$, the restriction $\mathrm{cok}:\mathrm{Ar}^{\rm im}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$ are quasi-inverse functors. $\Box$
For any pointed symmetric monoidal closed category $\mathcal{M}$ with a monoidal unit $V$, let $\mathrm{Alg}_{\mathcal{O}}(\mathcal{M})_{V//V}$ denote the full subcategory of $\mathrm{Alg}_{\mathcal{O}^{\otimes}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))$ spanned by augmentations of $\mathcal{O}^{\otimes}$-algebra objects.
Definition 5.5.
Let $\mathcal{M}$ be a locally presentable symmetric monoidal abelian category with a monoidal unit $V$. A commutative non-unital monoid object of $\mathcal{M}$ is a commutative Smith ideal $j:I\to R$ in the localized full subcategory $\mathrm{Ar}^{\rm im}(\mathcal{M})$ satisfying that the cokernel $\mathrm{cok}(j):R\to\mathrm{Coker}(j)$ is isomorphic with an argumentation $\varepsilon_{R}:R\to V$, and $\mathrm{Alg}_{\mathcal{O}}^{\rm nu}(\mathcal{M})$ denote the full subcategory of $\mathrm{Alg}_{\mathcal{O}^{\Box}}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ spanned by non-unital $\mathcal{O}^{\Box}$-algebra objects. In particular, $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ denotes the full subcategory $\mathrm{CAlg}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ spanned by commutative Smith ideals whose cokernels are isomorphic to $V$. We say that an object of $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ is a non-unital commutative $V$-algebra object of $\mathcal{M}$.
Proposition 5.6.
Let $\mathcal{M}$ be a locally presentable symmetric monoidal abelian category with a monoidal unit $V$. The categorical equivalences
$\mathrm{cok}:\mathrm{Ar}^{\rm im}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$
induces a categorical equivalence
$\mathrm{Alg}^{\rm nu}_{\mathcal{O}^{\Box}}(\mathcal{M})\to\mathrm{Alg}_{\mathcal{O}^{\otimes}}(\mathcal{M})_{V//V}.$
  • proof.
    By definition of the category $\mathrm{Alg}_{\mathcal{O}^{\Box}}^{\rm nu}(\mathcal{M})$, one has an isomorphism $V\oplus I\to R$. The assertion is clear. $\Box$
5.3. Definition of operadic non-unital algebra objects of symmetric monoidal categories
We generalize the categorical equivalence in Proposition 5.6. The main work of this section is to give a suitable definition of non-unital $\mathcal{O}$-algebra objects, which is Definition 5.7.
By the notation in Definition 5.2, we establish a definition of operadic non-unital algebra objects of symmetric monoidal model categories.
Definition 5.7.
Let $\mathcal{M}$ be a symmetric monoidal closed model category with a monoidal unit $V$. An homotopically augmented $\mathcal{O}$-algebra over $V$ $f:A\to B$ is an object of $\mathrm{Alg}_{\overrightarrow{\mathcal{O}^{\otimes}}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))$ that the unit $U^{0}(V)=(\mathrm{id}_{V}:V\to V)\to(f:A\to B)$ has a homotopy section. Let $\mathrm{Alg}_{\mathcal{O}^{\otimes}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))_{V//V}$ denotes the semi-model category of homotopically augmented $V$-algebras. In particular, $\mathrm{CAlg}(\mathcal{M})_{V//V}$ denotes the semi-model category of homotopically augmented $V$-algebras.
A non-unital $\mathcal{O}$-algebra over $V$ is a $\mathcal{O}$-Smith ideal whose cokernel is a homotopically augmented $\mathcal{O}$-algebra. Let $\mathrm{Alg}_{\mathcal{O}}^{\rm nu}(\mathcal{M})$ denote the full subcategory of $\mathrm{Alg}_{\mathcal{O}^{\Box}}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ spanned by non-unital $\mathcal{O}^{\Box}$-algebra objects. In particular, $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ denotes the semi-model category of non-unital commutative $V$-algebra object of $\mathcal{M}$.
Due to the following White and Yau’s theorem, we have a Quillen equivalence between semi-model categories of non-unital non-unital $\mathcal{O}$-algebras and homotopically augmented $\mathcal{O}$-algebras.
Theorem 5.8 ([WY24] Theorem 4.4.1).
Let $\mathcal{M}$ be a stable symmetric monoidal model category with a monoidal unit $V$. Assume that the adjunction $\mathrm{cok}:\mathrm{Ar}^{\Box}(\mathcal{M})\rightleftarrows\mathrm{Ar}^{\otimes}(\mathcal{M}):\mathrm{ker}$ induces a Quillen adjunction
$\mathrm{cok}:\mathrm{Alg}_{\mathcal{O}^{\Box}}(\mathrm{Ar}^{\Box}(\mathcal{M}))\rightleftarrows\mathrm{Alg}_{\mathcal{O}^{\otimes}}(\mathrm{Ar}^{\otimes}(\mathcal{M})):\mathrm{ker}$
between semi-model categories, and the colored operad $\mathcal{O}$ in $\mathcal{M}$ is cofibrant in $\mathrm{Ar}^{\Box}(\mathcal{M})^{\mathfrak{C}}$. Then the Quillen adjunction is a Quillen equivalence. $\Box$
Proposition 5.9.
Under the assumption of Theorem 5.8, the Quillen equivalence in Theorem 5.8 induces a Quillen equivalence
$\mathrm{cok}:\mathrm{Alg}_{\mathcal{O}^{\Box}}^{\rm nu}(\mathrm{Ar}^{\Box}(\mathcal{M}))\rightleftarrows\mathrm{Alg}_{\mathcal{O}^{\otimes}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))_{V//V}:\mathrm{ker}.$
$\Box$
White introduced the commutative monoid axiom of monoidal model categories.
Definition 5.10 ([Whi22], Definition 3.1 and 3.4).
A symmetric monoidal model category $\mathcal{M}$ is said to be satisfying the commutative monoid axiom if, for all $n>0$, the $n$-fold symmetric pushout product functor $(-)^{\Box n}/\Sigma_{n}:\mathcal{M}\to\mathcal{M}$ preserves all trivial cofibrations. Further if the $n$-fold symmetric pushout product functor preserves all cofibrations and trivial cofibrations, $\mathcal{M}$ is said to be satisfying the strong commutative monoid axiom.
In particular, by [WY24, Corollary 5.1.2], we obtain the following:
Corollary 5.11.
Let $\mathcal{M}$ be a cofibrantly generated stable model category satisfying the strong commutative monoid axiom, the monoid axiom, and in which cofibrant commutative $\mathrm{Ar}^{\Box}(\mathcal{M})$. Then the cokernel-and-kernel adjunction induces a Quillen equivalence
$\mathrm{cok}:\mathrm{CAlg}^{\rm nu}(\mathcal{M})\rightleftarrows\mathrm{CAlg}(\mathcal{M})_{V//V}:\mathrm{ker}$
between model categories. $\Box$
Remark 5.12.
If $\mathcal{M}$ is a locally presentable stable symmetric monoidal model category, the presentable $\infty$-category of augmented algebra is represented by the model category $\mathrm{CAlg}(\mathcal{M})_{V//V}$. Hence the left-hand-side $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ is equivalent to the $\infty$-category defined by Lurie [Lur17, p.949, Definition 5.4.4.9 and Proposition 5.4.4.10].

### v5

@@
 Abstract.
-This article is a generalization of a result in Quillen’s note [Qui96] “Module theory over non-unital rings” giving a one-to-one correspondence between bilocalization of abelian categories of modules and idempotent ideals of the base ring.
-Faltings [Fal88]; Gabber and Ramero [GR03] established almost mathematics, the same as Quillen’s bilocalization of a category of modules by nil-modules.
-In this paper, by using the theory of Smith ideals mentioned in Hovey [Hov14], we consider almost mathematics of symmetric monoidal pointed model categories and prove a weak analogue of the one-to-one correspondence in Quillen’s note [Qui96].
+This article is a generalization of a result in Quillen’s note [14] “Module theory over non-unital rings” giving a one-to-one correspondence between bilocalization of abelian categories of modules and idempotent ideals of the base ring.
+Faltings [4]; Gabber and Ramero [5] established almost mathematics, the same as Quillen’s bilocalization of a category of modules by nil-modules.
+In this paper, by using the theory of Smith ideals mentioned in Hovey [9], we consider almost mathematics of symmetric monoidal pointed model categories and prove a weak analogue of the one-to-one correspondence in Quillen’s note [14].
 Key words and phrases:
-Symmetric monoidal model categories, Smith ideals, almost mathematics, Bousfield localization and colocalization
+Almost mathematics, Smith ideals, monoidal model categories, Bousfield localization and colocalization
+1991 Mathematics Subject Classification:
+18N55 (primary), 18N70 (secondary)
+The author was supported by Grants-in-Aid for Scientific Research No. 23K030080, Japan Society for the Promotion of Science.
 1. Introduction
-Faltings [Fal88] first introduced one of the topics of algebra theory almost mathematics.
-Later, Gabber and Ramero provided a textbook [GR03] of detailed foundation of almost mathematics, where they called those materials of the theory almost ring theory, almost modules, almost algebras, and almost homotopy algebras.
-Referring Gabber and Ramero’s book, almost mathematics is obtained by a localization of an abelian category of modules by almost isomorphisms: Let $V$ be a unital commutative ring with an idempotent ideal $\mathcal{m}$ of $V$.
-A $V$-module $M$ is said to be almost zero if $M$ is killed by $\mathcal{m}$, and almost mathematics is a theory of algebra working by localizing the category of $V$-modules by the Serre subcategory of almost zero modules.
+Faltings [4] first introduced the theory of almost mathematics to apply $p$-adic Hodge theory, and subsequently, Gabber and Ramero [5] developed the theory of algebra within this framework.
+In their terminology, the theory includes almost ring theory, almost modules, almost algebras, and almost homotopy algebras.
+Building on Gabber and Ramero’s work, almost mathematics is formulated by localizing an abelian category of modules at almost isomorphisms: let $V$ be a unital commutative ring with an idempotent ideal $\mathfrak{m}$ of $V$.
+A $V$-module $M$ is defined to be almost zero if it is annihilated by $\mathfrak{m}$, and almost mathematics is an algebraic theory constructed by localizing the category of $V$-modules with respect to the Serre subcategory of almost zero modules.
 Almost isomorphisms are morphisms whose kernels and cokernels are both almost zero.
-In the localized category, module objects are called almost modules and algebra objects are called almost algebras.
-Independently, Quillen considered the same theory of almost mathematics in unpublished note [Qui96] as linear algebra over non-unital rings, which he characterized as bilocalization - both localization and colocalization - of abelian categories of modules.
-Further, Quillen proved the following one-one correspondence between idempotent ideals and reflexive bilocalization functors.
+In the localized category, module objects are referred to as almost modules , while algebra objects are called almost algebras.
+Independently, in an unpublished note [14], Quillen considered the same theory of almost mathematics as linear algebra over non-unital rings.
+He characterized it as a bilocalization, that is, both a localization and a colocalization, of abelian categories of modules.
+Further, Quillen proved the following one-to-one correspondence between idempotent ideals and reflexive bilocalization functors.
 Theorem 1.1.
-([Qui96] Proposition 6.5) Let $V$ be a commutative ring with a multiplicative unit.
-There is a one-to-one correspondence between Serre subcategories $\mathcal{S}$ of $\mathrm{Mod}_{V}$ that the localization $F:\mathrm{Mod}_{V}\to\mathrm{Mod}_{V}/\mathcal{S}$ is also a colocalization, and idempotent ideals of $V$.
+([14] Proposition 6.5) Let $V$ be a commutative ring with a multiplicative unit.
+There is a one-to-one correspondence between idempotent ideals of $V$ and Serre subcategories $\mathcal{S}$ of $\mathrm{Mod}_{V}$ for which the localization $F:\mathrm{Mod}_{V}\to\mathrm{Mod}_{V}/\mathcal{S}$ is also a colocalization.
 $\Box$
-The purpose of this article is to establish an almost mathematics theory of symmetric monoidal model categories, where we prove a monoidal categorical analogue of Theorem 1.1.
-Namely, by utilizing the commutative Smith ideal theory developed by Hovey [Hov14] and White–Yau [WY24], we introduce homotopy analogues of idempotent ideals and almost isomorphisms, which we refer to as homotopically idempotent Smith ideals and almost weak equivalences.
-The main results of this paper are as follows, as mentioned in Theorem 2.11.
+The purpose of this article is to establish a theory of almost mathematics for symmetric monoidal model categories and to prove a monoidal categorical analogue of Theorem 1.1.
+Namely, using the commutative Smith ideal theory developed by Hovey [9] and White–Yau [18], we introduce homotopy analogues of idempotent ideals and almost isomorphisms, which we refer to as homotopically idempotent Smith ideals and almost weak equivalences.
+The main result of this paper are summarized in Theorem 2.12.
 Theorem A.
-Let $\mathcal{M}$ be a pointed symmetric monoidal model category with a monoidal unit object $V$, $j:I\to V$ be a homotopically idempotent commutative Smith ideal. Then the internal-hom functor
-$\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-):\mathcal{M}\to\mathcal{M}$
-is equivalent to the Bousfield bilocalization functor by almost weak equivalences.
-Since $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ is a right Quillen functor, this theorem shows us that the Bousfield localization by almost weak equivalences is both a left and right Quillen functor.
-Conversely, in the case that the monoidal unit generates the stable symmetric monoidal model category in the sense of MacLane [Mac88, Chapter V, Section 7, p.127], we can construct a homotopically idempotent ideal (Proposition 4.5 and Theorem 4.7).
-Furthermore, we generalize some result of almost algebras for model category setting by using Smith ideal language.
+Let $\mathcal{M}$ be a pointed symmetric monoidal model category with a monoidal unit object $V$, and let $j:I\to V$ be a homotopically idempotent commutative Smith ideal.
+Then the internal-hom functor
+[…]
+is equivalent to the Bousfield bilocalization functor with respect to almost weak equivalences.
+[…]
+Conversely, in the case that the monoidal unit generates the stable symmetric monoidal model category in the sense of MacLane [11, Chapter V, Section 7, p.127], we can construct a homotopically idempotent ideal (Proposition 4.5 and Theorem 4.7).
+Furthermore, we generalize some results on almost algebras to the model category setting using the language of Smith ideals.
 Again, let $V$ be a commutative ring with an idempotent ideal $\mathfrak{m}$.
@@
 $A_{!!}=V\amalg_{\tilde{\mathfrak{m}}}(\tilde{\mathfrak{m}}\otimes_{V}A)$
-for any almost unital $V$-algebra $A$ in [GR03, Section 2.2].
-Those morphisms $\tilde{\mathfrak{m}}\to V\to A$ and $\tilde{\mathfrak{m}}\to\tilde{\mathfrak{m}}\otimes_{V}A\to A$ induce a ring homomorphism $A_{!!}\to A$, which is almost isomorphism.
-Due to [GR03, Proposition 2.2.29], the almost unitalization $(-)_{!!}$ is a left adjoint of the localization $\mathrm{CAlg}(\mathrm{Mod}_{V})\to\mathrm{CAlg}(\mathrm{Mod}_{V}^{\rm al})$.
+for any almost unital $V$-algebra AA in [5, Section 2.2].
+Those morphisms $\tilde{\mathfrak{m}}\to V\to A$ and $\tilde{\mathfrak{m}}\to\tilde{\mathfrak{m}}\otimes_{V}A\to A$ induce a ring homomorphism $A_{!!}\to A$, which is an almost isomorphism.
+Due to [5, Proposition 2.2.29], the almost unitalization $(-)_{!!}$ is a left adjoint of the localization $\mathrm{CAlg}(\mathrm{Mod}_{V})\to\mathrm{CAlg}(\mathrm{Mod}_{V}^{\rm al})$.
 We call the left adjoint $(-)_{!!}$ almost unitalization in this paper.
-By using Smith ideal theory, we generalize the almost unitalization functor $(-)_{!!}$ that is right adjoint of the localization functor.
+Using Smith ideal theory, we generalize the almost unitalization functor $(-)_{!!}$ in the language of Smith ideals.
 In Section 2, we introduce the theory of almost mathematics of pointed symmetric monoidal model categories; homotopically idempotent commutative Smith ideals, homotopically almost zero objects, and homotopically almost weak equivalences.
-In Section 3, we generalize the above almost unitalization functor by Smith ideal theory language, and we prove that it is a left adjoint of the Bousfield localization functor by the class of almost weak equivalences defined in Section 2, as Theorem 3.6 and Corollary 3.7.
-In Section 4, as the inverse direction of Theorem 2.11, we demonstrate constructing homotopically idempotent Smith ideal from given symmetric monoidal bilocalization functor on symmetric monoidal model categories under some mild conditions, and the result is proven as Theorem 4.7.
-Finally, in Section 5, we remark that the cokernel-and-kernel functors in the (operadic) Smith ideal theory induces a categorical equivalence between non-unital algebra objects and augmented algebra objects of stable symmetric monoidal model categories.
+In Section 3, we generalize the above almost unitalization functor in the language of Smith ideal theory, and we prove that it is a left adjoint of the Bousfield localization functor with respect to the class of almost weak equivalences defined in Section 2; see Theorem 3.6 and Corollary 3.7.
+In Section 4, as the inverse direction of Theorem 2.12, we construct a homotopically idempotent Smith ideal from a given symmetric monoidal bilocalization functor on a symmetric monoidal model category under mild conditions.
+The result is proved as Theorem 4.7.
+Finally, in Section 5, we remark that the cokernel-and-kernel functors in the (operadic) Smith ideal theory induce a categorical equivalence between non-unital algebra objects and augmented algebra objects of stable symmetric monoidal model categories.
+For the reader’s convenience, we mention related independent work of Hebestreit and Scholze on higher almost ring theory [7].
+Their work develops a higher almost mathematics framework in an $\infty$-categorical setting.
+The present paper is complementary, developing a model-categorical framework based on Smith ideal techniques.
+To clarify the independent development of the two projects, we also record that the fourth version of the present preprint appeared as arXiv:2301.04541v4 on July 29, 2024 UTC (July 30 in Japan), and that the first version of Hebestreit and Scholze’s preprint appeared as arXiv:2409.01940v1 on September 3, 2024.
 1.1. Example: The almost mathematics of a stable model category of chain complexes
-In this section, we consider the model category of chain complexes of modules as an example of almost mathematics of model categories.
-We fix a commutative base ring $V$ and an idempotent ideal $\mathfrak{m}$ ($\mathfrak{m}^{2}=\mathfrak{m}$) of $V$, and we assume that the tensor product $\tilde{\mathfrak{m}}=\mathfrak{m}\otimes_{V}\mathfrak{m}$ is a flat $V$-module.
-Let $\operatorname{Ch}(V)$ denote the category of chain complexes of $V$-modules.
-Then $\operatorname{Ch}(V)$ has a canonical model structure called the injective model structure that cofibrations are degree-wise injections and weak equivalences are quasi-isomorphisms.
-Since we assume that $\tilde{\mathfrak{m}}$ is a flat $V$-module, the functor $\tilde{\mathfrak{m}}\otimes_{V}(-):\operatorname{Ch}(V)\to\operatorname{Ch}(V)$ is exact and a left Quillen functor, whose right adjoint is the mapping complex functor $\mathrm{Map}_{V}(\tilde{\mathfrak{m}},\,-)$.
+[…]
 Example 1.2.
-A perfectoid commutative valuation ring is defined to be a valuation ring of a perfectoid field defined in Scholze’s paper [Sch12, p.246, Definition 1.2].
+A perfectoid commutative valuation ring is a valuation ring of a perfectoid field in the sense of Scholze [15, p.246, Definition 1.2].
 Let $V$ be a perfectoid valuation ring.
-Then the set of topologically nilpotent elements $\mathfrak{m}$ of $V$ is an idempotent maximal ideal and a flat $V$-module, explained in, for example, Bhatt’s lecture note [Bha17, p.18, Example 4.1.3].
+Then the set of topologically nilpotent elements $\mathfrak{m}$ of $V$ is an idempotent maximal ideal and a flat $V$-module, as explained, for example, in Bhatt’s lecture note [3, p.18, Example 4.1.3].
 In general, the product map $\tilde{\mathfrak{m}}\to\mathfrak{m}$ is an isomorphism if $\mathfrak{m}$ is a flat $V$-module.
-In the category $\mathrm{Ch}(V)$ of almost weak equivalences is defined as the following:
+In the category $\mathrm{Ch}(V)$, almost weak equivalences are defined as follows:
 Definition 1.3.
 A morphism $f:M\to N$ of $V$-complexes is said to be an almost quasi-isomorphism if the induced morphism $\tilde{\mathfrak{m}}\otimes_{V}f:\tilde{\mathfrak{m}}\otimes_{V}M\to\tilde{\mathfrak{m}}\otimes_{V}N$ is a quasi-isomorphism.
-Let $\operatorname{Ch}(V)^{\rm al}$ denote the Bousfield localization of $\operatorname{Ch}(V)$ by the class of almost quasi-isomorphisms.
+Let $\operatorname{Ch}(V)^{\rm al}$ denote the Bousfield localization of the model category $\operatorname{Ch}(V)$ by the class of almost quasi-isomorphisms.
 A chain complex $E$ is said to be almost local if $E$ is a local object with respect to almost quasi-isomorphisms.
@@
 is an isomorphism in the derived category $D(V)$ of $\operatorname{Ch}(V)$.
-In Quillen’s note [Qui96], a $V$-module $E$ that the induced homomorphism $E\to\mathrm{Hom}_{V}(\mathfrak{m},\,E)$ is an isomorphism is said to be closed, and closed modules are almost local objects and vice versa.
+In Quillen’s note [14], a $V$-module $E$ is said to be closed if the induced homomorphism $E\to\mathrm{Hom}_{V}(\mathfrak{m},\,E)$ is an isomorphism, and it is shown that closed modules are almost local objects and vice versa; see [14, Proposition 5.3].
 The following lemma characterizes almost local objects.
@@
 $E\to\mathrm{Map}_{V}(\tilde{\mathfrak{m}},\,E).$
-  • proof.
 The only if direction is clear by definition of almost local objects.
 Conversely, assume that $E$ is almost local.
-Since the canonical morphism $\tilde{\mathfrak{m}}\otimes_{V}M\to M$ is an almost quasi-isomorphism for any $M\in\operatorname{Ch}(V)$, then the induced morphism
+Since the canonical morphism $\tilde{\mathfrak{m}}\otimes_{V}M\to M$ is an almost quasi-isomorphism for any $M\in\operatorname{Ch}(V)$, the induced morphism
 $\mathrm{Hom}_{D(V)}(M,\,E)\to\mathrm{Hom}_{D(V)}(M,\,\mathrm{Map}_{V}(\tilde{\mathfrak{m}},\,E))\simeq\mathrm{Hom}_{D(V)}(\tilde{\mathfrak{m}}\otimes_{V}M,\,E).$
@@
 Theorem 1.5.
-Then the mapping complex functor
+The mapping complex functor
 $\mathrm{Map}_{V}(\tilde{\mathfrak{m}},\,-):\operatorname{Ch}(V)\to\operatorname{Ch}(V)^{\rm al}$
 is equivalent to the localization functor $L^{\rm al}:\operatorname{Ch}(V)\to\operatorname{Ch}(V)^{\rm al}$ on the derived categories.
-  • proof.
 This is the special case of Theorem 2.12. $\Box$
-Acknowledgments
-I would like to thank anonymous referees for giving me innovative ideas and constructive comments to use Hovey’s Smith ideal theory and White and Yau’s operadic Smith ideal theory.
 2. Almost mathematics of pointed symmetric monoidal model categories
-In this section, referring White and Yau’s work [WY24], we recall definition and some properties of symmetric monoidal model model categories and (commutative) Smith ideals of them.
-Next, in order to establish almost mathematics, we define homotopically idempotent Smith ideals, and we prove this paper’s main result Theorem 2.11.
-[…]
+In this section, following White and Yau’s work [18], we recall the definition and some properties of symmetric monoidal model categories and their (commutative) Smith ideals.
+Next, in order to establish almost mathematics, we define homotopically idempotent Smith ideals, and we prove the main result, Theorem 2.12.
+A category is pointed if it has an initial object $\emptyset$ and a final object * such that the unique morphism $\emptyset\to*$ is an isomorphism.
+We call the initial object the zero object, and 0 denotes it.
+Throughout this paper, we consider pointed (model) categories.
 Let $\mathcal{M}$ be a model category.
@@
 It is well-known that the following adjunctions
-${\mathcal{M}^{\rm cof}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R}$$\scriptstyle{Q}$${\mathcal{M}^{\rm fib}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
+[…]
 are Quillen equivalences, where $\mathcal{M}^{\rm cof}$ (resp. $\mathcal{M}^{\rm fib}$) denotes the full subcategory spanned by all cofibrant (resp. fibrant) objects.
 2.1. Commutative Smith ideals of symmetric monoidal model categories
-A symmetric monoidal model category $\mathcal{M}$ is a model category with a symmetric monoidal structure $-\otimes-:\mathcal{M}\times\mathcal{M}\to\mathcal{M}$ satisfying the following two axioms:
+A symmetric monoidal model category $\mathcal{M}$ is a model category with a symmetric monoidal structure $\otimes-:\mathcal{M}\times\mathcal{M}\to\mathcal{M}$ satisfying the following two axioms:
   1. (1)
@@
 Further, if $f$ or $g$ is a trivial cofibration, so is $f\Box g$.
-We say that the symmetric monoidal structure on $\mathcal{M}$ is closed if $\mathcal{M}$ contains an internal hom-object $\mathrm{Map}_{\mathcal{M}}(M,\,N)$ with respect to the monoidal structure $-\otimes-:\mathcal{M}\times\mathcal{M}\to\mathcal{M}$ for any $M,\,N\in\mathcal{M}$.
-In this case, the push-out product axiom implies that, for any cofibrant object $M$, both those induced functors $-\otimes M:\mathcal{M}\to\mathcal{M}$ and $M\otimes-:\mathcal{M}\to\mathcal{M}$ are left Quillen functors.
+We say that the symmetric monoidal structure on $\mathcal{M}$ is closed if $\mathcal{M}$ contains an internal hom-object $\mathrm{Map}_{\mathcal{M}}(M,\,N)$ with respect to the monoidal structure $\otimes-:\mathcal{M}\times\mathcal{M}\to\mathcal{M}$ for any $M,\,N\in\mathcal{M}$.
+In this case, the push-out product axiom implies that, for any cofibrant object $M$, both induced functors $\otimes M:\mathcal{M}\to\mathcal{M}$ and $M\otimes-:\mathcal{M}\to\mathcal{M}$ are left Quillen functors.
 Further, we assume that all model categories in this paper are cofibrantly generated.
@@
 Further, if a cofibrantly generated model category is locally presentable, it is said to be combinatorial.
-By the theorem of Barwick and Smith [Bar10, Theorem 4.7], for any left proper and combinatorial model category and a small set of homotopy class of morphisms of it, the Bousfield localization exists such that the localized model category is also left proper and combinatorial.
-Let $\Delta^{1}$ denote the category with two objects 00 and $1$, and only one non-trivial morphism $0\to 1$.
-The functor category $\mathrm{Fun}(\Delta^{1},\,\mathcal{M})$ from $\Delta^{1}$ to $\mathcal{M}$ is called the arrow category of $\mathcal{M}$, and $\mathrm{Ar}(\mathcal{M})$ denotes it.
-By Hovey [Hov14, Theorem 1.2], the arrow category $\mathrm{Ar}(\mathcal{M})$ has two distinct symmetric monoidal structures derived from $\mathcal{M}$’s: For any two morphisms in $\mathcal{M}$, $f:X_{0}\to X_{1}$ and $g:Y_{0}\to Y_{1}$, one has a commutative square:
+By the theorem of Barwick and Smith [2, Theorem 4.7], for any left proper and combinatorial model category and a small set of homotopy classes of morphisms in it, the Bousfield localization exists such that the localized model category is also left proper and combinatorial.
+Let $\Delta^{1}$ denote the category with two objects 0 and $1$, and only one non-trivial morphism $0\to 1$.
+The functor category $\mathrm{Fun}(\Delta^{1},\,\mathcal{M})$ from $\Delta^{1}$ to $\mathcal{M}$ is called the arrow category of $\mathcal{M}$ and is denoted by $\mathrm{Ar}(\mathcal{M})$.
+By Hovey [9, Theorem 1.2], the arrow category $\mathrm{Ar}(\mathcal{M})$ has two distinct symmetric monoidal structures derived from that of $\mathcal{M}$.
+For any two morphisms $f:X_{0}\to X_{1}$ and $g:Y_{0}\to Y_{1}$ in $\mathcal{M}$, one has a commutative square:
 ${X_{0}\otimes Y_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f\otimes\mathrm{id}}$$\scriptstyle{\mathrm{id}\otimes g}$${X_{1}\otimes Y_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{id}\otimes g}$${X_{0}\otimes Y_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f\otimes\mathrm{id}}$${X_{1}\otimes Y_{1}.}$
-One of them is the diagonal (or tensor) product monoidal structure is defined by $f\otimes g$ as the composition
+The diagonal (or tensor) product monoidal structure is defined by $f\otimes g$ as the composition
 $f\otimes g=(f\otimes\mathrm{id})\circ(\mathrm{id}\otimes g)=(\mathrm{id}\otimes g)\circ(f\otimes\mathrm{id}):X_{0}\otimes Y_{0}\to X_{1}\otimes Y_{1}.$
@@
 $f\Box g:(X_{0}\otimes Y_{1})\amalg_{X_{0}\otimes Y_{0}}(X_{1}\otimes Y_{0})\to X_{1}\otimes Y_{1}.$
-[…]
+The arrow category $\mathrm{Ar}(\mathcal{M})$ has the evaluation functors $\mathrm{Ev}_{i}(f:X_{0}\to X_{1})=X_{i}\ (i=0,\,1)$, which have a left adjoint and a right adjoint:
 ${\mathrm{Ev}_{i}:\mathrm{Ar}(\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\mathcal{M}:L_{i},\,U_{i}\ (i=0,\,1).}$
@@
 degree-wise cofibrations).
-Those monoidal model categories $\mathrm{Ar}^{\Box}(\mathcal{M})$ and $\mathrm{Ar}^{\otimes}(\mathcal{M})$ are connected the cokernel-and-kernel adjunction
+These monoidal model categories $\mathrm{Ar}^{\Box}(\mathcal{M})$ and $\mathrm{Ar}^{\otimes}(\mathcal{M})$ are connected by the cokernel-and-kernel adjunction
 $\mathrm{cok}:\mathrm{Ar}^{\Box}(\mathcal{M})\rightleftarrows\mathrm{Ar}^{\otimes}(\mathcal{M}):\mathrm{ker},$
-which is a Quillen adjunction. The cokernel functor $\mathrm{cok}$ is defined by the sending
+which is a Quillen adjunction. The cokernel functor $\mathrm{cok}$ is defined by
 $\mathrm{cok}:\mathrm{Ar}^{\Box}(\mathcal{M})$ $\to\mathrm{Ar}^{\otimes}(\mathcal{M})$
@@
 respectively.
-In the case that $\mathcal{M}$ is a pointed, left proper, and combinatorial symmetric monoidal model category, the adjunction is lax monoidal Quillen by Hovey’s work (the proof is available at [WY24, Proposition 2.4.3]).
+In the case that $\mathcal{M}$ is a pointed, left proper, and combinatorial symmetric monoidal model category, the adjunction is lax monoidal Quillen by Hovey’s work (the proof is available at [18, Proposition 2.4.3]).
 Further, if $\mathcal{M}$ is stable, in addition, the adjunction $(\mathrm{cok},\,\mathrm{ker})$ is a Quillen equivalence.
-(See Hovey [Hov14, Theorem 1.4] or White and Yau [WY24, Proposition 2.3]).
+(See Hovey [9, Theorem 1.4] or White and Yau [18, Proposition 2.3]).
 We recall the definition of commutative Smith ideals.
-A commutative monoid object $A$ of $\mathcal{M}$ is a monoid whose product $\mu_{A}:A\otimes A\to A$ commutes the braiding $\tau_{A,\,A}:A\otimes A\to A\otimes A$, denoting $\mathrm{CAlg}(\mathcal{M})$ the subcategory of commutative objects of $\mathcal{M}$.
-Definition 2.1 (c.f. [WY24] Definition 3.3.1).
+A commutative monoid object $A$ of $\mathcal{M}$ is a monoid whose product $\mu_{A}:A\otimes A\to A$ commutes with the braiding $\tau_{A,\,A}:A\otimes A\to A\otimes A$.
+We denote by $\mathrm{CAlg}(\mathcal{M})$ the subcategory of commutative objects of $\mathcal{M}$.
+Definition 2.1 (c.f. [18] Definition 3.3.1).
 Let $\mathcal{M}$ be a pointed symmetric monoidal model category. A commutative Smith ideal is a commutative monoid object of the symmetric monoidal model category $\mathrm{Ar}^{\Box}(\mathcal{M})$.
@@
   • (2)
-    Write $\tilde{j}:I\otimes I\to I\otimes V\to V$ and $\tilde{I}=I\otimes I$.
-Then $\tilde{I}$ is a cofibrant object and $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ has a homotopy section $\Delta:\tilde{I}\to\tilde{I}\otimes\tilde{I}$ satisfying the associativity:
+    Write $\tilde{I}=I\otimes I$.
+Then $\tilde{I}$ is cofibrant and $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ has a homotopy section $\Delta:\tilde{I}\to\tilde{I}\otimes\tilde{I}$ satisfying the associativity:
 […]
   • (3)
-    The tensor product $\tilde{I}\otimes\tilde{I}$ is a group like object.
-That is, the unit morphism $\tilde{I}\otimes\tilde{I}\to\Omega\Sigma(\tilde{I}\otimes\tilde{I})$ is a weak equivalence, where $\Omega$ is the pointed loop functor and $\Sigma$ the suspension functor on the pointed model category $\mathcal{M}$.
+[…]
 Proposition 2.3.
-Let $\mathcal{M}$ be a symmetric monoidal model category with a monoidal unit $V$, and $j:\tilde{I}\to V$ a homotopically idempotent commutative Smith ideal of $V$.
+Let $\mathcal{M}$ be a symmetric monoidal model category with a monoidal unit $V$, and $j:I\to V$ a homotopically idempotent commutative Smith ideal of $V$.
 Then the product $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ is a weak equivalence.
-  • proof.
 We show that the homotopy kernel of $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ is contractible. Let $K$ denote the homotopy kernel of $\mu_{\tilde{I}}$. One has a diagram
@@
 is an isomorphism, satisfying the coherence condition in Definition 2.2.
+Example 2.6 ([6, 12] Torsion and derived completion).
+Let $A$ be a commutative ring and $\mathfrak{a}\subset A$ a finitely generated weakly proregular ideal.
+Consider the injective model category $\operatorname{Ch}(A)$ and let $K_{\infty}(\mathfrak{a})$ denote the infinite dual Koszul complex.
+Then, on the derived category $D(A)$, the derived torsion functor $\mathbb{R}\Gamma_{\mathfrak{a}}(-)$ is equivalent to the tensor functor $K_{\infty}(\mathfrak{a})\otimes_{A}^{\mathbb{L}}(-)$, and the derived completion functor $\mathbb{L}\Lambda_{\mathfrak{a}}(-)$ is equivalent to the internal-hom functor $\mathrm{Map}_{A}(K_{\infty}(\mathfrak{a}),\,-)$.
+Note that $K_{\infty}(\mathfrak{a})$ is homotopy idempotent: $K_{\infty}(\mathfrak{a})\otimes_{A}^{\mathbb{L}}K_{\infty}(\mathfrak{a})\simeq K_{\infty}(\mathfrak{a})$.
+Therefore, the adjunction
+$K_{\infty}(\mathfrak{a})\otimes_{A}^{\mathbb{L}}(-):D(A)\rightleftarrows D(A):\mathrm{Map}_{A}(K_{\infty}(\mathfrak{a}),\,-)$
+gives a bilocalization of $D(A)$, where derived torsion is realized as the tensor side and derived completion as the internal-hom side.
+In this sense, the derived completion functor is an instance of the internal-hom side of our general tensor–internal-hom bilocalization.
+This viewpoint is compatible with the equivalence between firm and closed modules over an idempotent ideal: derived $\mathfrak{a}$-torsion complexes are the fixed objects of $K_{\infty}(\mathfrak{a})\otimes_{A}^{\mathbb{L}}(-)$, while derived $\mathfrak{a}$-complete complexes are the fixed objects of $\mathrm{Map}_{A}(K_{\infty}(\mathfrak{a}),\,-)$.
+The adjunction restricts to the Greenlees–May duality between the corresponding full subcategories [12, 13].
+Hence the commutative Smith ideal $j:I\to A$ obtained from a cofibrant replacement of $K_{\infty}(\mathfrak{a})\to A$ gives a fundamental example of a homotopically idempotent commutative Smith ideal governing this duality.
 Definition 2.7.
 Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal of the monoidal unit $V$.
-A morphism $f:M\to N$ in $\mathcal{M}$ is an almost weak equivalence if the induced morphism $\tilde{I}\otimes Q(f):\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(N)$ is a weak equivalence, and $\mathcal{M}^{\rm al}$ denotes the Bousfield localization of $\mathcal{M}$ by the class of almost weak equivalences.
+A morphism $f:M\to N$ in $\mathcal{M}$ is called an almost weak equivalence if the induced morphism $\tilde{I}\otimes Q(f):\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(N)$ is a weak equivalence, and $\mathcal{M}^{\rm al}$ denotes the Bousfield localization of $\mathcal{M}$ with respect to the class of almost weak equivalences.
 Definition 2.8.
-An object $M$ of $\mathcal{M}$ is almost local if the induced morphism $f^{*}:\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(N_{2},\,M)\to\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(N_{1},\,M)$ is an isomorphism.
-Dually, $M$ is almost colocal if, for any homotopically almost weak equivalence $f:N_{1}\to N_{2}$, the induced map $f_{*}:\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(M,\,N_{1})\to\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(M,\,N_{2})$ is an isomorphism.
+An object $M$ of $\mathcal{M}$ is almost local if, for any almost weak equivalence $f:N_{1}\to N_{2}$, the induced morphism
+$f^{*}:\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(N_{2},\,M)\to\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(N_{1},\,M)$
+is an isomorphism. Dually, $M$ is almost colocal if, for any almost weak equivalence $f:N_{1}\to N_{2}$, the induced map
+$f_{*}:\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(M,\,N_{1})\to\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(M,\,N_{2})$
+is an isomorphism.
 Definition 2.9.
 Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal. An object $E$ is homotopically $I$-closed if the induced morphism
-$\mathrm{Map}_{\mathcal{M}}(\tilde{j},\,E):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,E)\to\mathrm{Map}_{\mathcal{M}}(V,\,E)\to E$
+$\mathrm{Map}_{\mathcal{M}}(\tilde{j},\,E):\mathrm{Map}_{\mathcal{M}}(V,\,E)\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,E)$
 is a weak equivalence. An object $E$ is homotopically $I$-firm if $\tilde{j}:\tilde{I}\to V$ induces a weak equivalence
@@
 Remark 2.10.
-In Quillen’s note [Qui96, Definition 5.2], $I$-closed modules said to be closed.
-However, this paper considers closed monoidal model categories, so we use the terminology $I$-closed objects for given a homotopically idempotent Smith ideal $I$.
+In Quillen’s note [14, Definition 5.2], $I$-closed modules are said to be closed.
+However, this paper considers closed monoidal model categories, so we use the terminology $I$-closed objects for a given homotopically idempotent Smith ideal $I$.
 We verify the equivalence between almost local objects and homotopically $I$-closed objects.
@@
   1. (1)
-    For any object $M$, for any object $M$, $\tilde{j}:\tilde{I}\to V$ induces an almost weak equivalence $\tilde{j}:\tilde{I}\otimes M\to V\otimes M\to M$.
+    For any object $M$, $\tilde{j}:\tilde{I}\to V$ induces an almost weak equivalence $\tilde{j}:\tilde{I}\otimes M\to V\otimes M\to M$.
   2. (2)
     An object $E$ is almost local if and only if it is homotopically $I$-closed.
-  • proof.
 To prove (1), consider $Q(\tilde{j}):\tilde{I}\to Q(V)\to V$ as a factorization, where the first morphism is a cofibration and the second is a trivial fibration.
 Then, by the unit condition of symmetric monoidal model categories, $\tilde{I}\otimes Q(V)\otimes Q(M)\to\tilde{I}\otimes V\otimes Q(M)$ is a weak equivalence by definition.
-Additionally, since $\tilde{I}\otimes\tilde{I}\to\tilde{I}\otimes Q(V)$ is a weak equivalence of cofibrant objects, and $-\otimes Q(M)$ is a left Quillen functor, by Ken Brown’s lemma [Hov99, p.6 Lemma 1.1.12], we have that $\tilde{I}\otimes\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(V)\otimes Q(M)$ is also a weak equivalence.
+Additionally, since $\tilde{I}\otimes\tilde{I}\to\tilde{I}\otimes Q(V)$ is a weak equivalence of cofibrant objects, and $\otimes Q(M)$ is a left Quillen functor, by Ken Brown’s lemma [8, p.6 Lemma 1.1.12], we have that $\tilde{I}\otimes\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(V)\otimes Q(M)$ is also a weak equivalence.
 Consequently, $\tilde{I}\otimes\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(M)$ is a weak equivalence.
@@
 Theorem 2.12.
-Let $\mathcal{M}$ be a pointed symmetric monoidal model category with a monoidal unit object $V$, $j:I\to V$ be a homotopically idempotent commutative Smith ideal and $\mathcal{M}^{\rm al}$ denoting the Bousfield localization by the collection of almost weak equivalences.
-Then, the internal-hom functor
+Let $\mathcal{M}$ be a pointed symmetric monoidal model category with a monoidal unit object $V$, let $j:I\to V$ be a homotopically idempotent commutative Smith ideal, and let $\mathcal{M}^{\rm al}$ denote the Bousfield localization with respect to the collection of almost weak equivalences.
+Then the internal-hom functor
 $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-):\mathcal{M}\to\mathcal{M}^{\rm al}$
 is equivalent to the Bousfield bilocalization functor.
-Furthermore, let $\mathcal{M}^{\rm fi}$ denote the homotopically essential image of $\tilde{I}\otimes Q:\mathcal{M}\to\mathcal{M}$. Then, the induced adjunction
-$\tilde{I}\otimes Q:\mathcal{M}^{\rm al}\to\mathcal{M}^{\rm fi}:\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$
+Furthermore, let $\mathcal{M}^{\rm fi}$ denote the homotopically essential image of $\tilde{I}\otimes Q:\mathcal{M}\to\mathcal{M}$. Then the induced adjunction
+[…]
 is a Quillen equivalence. In particular, for any object $E$, the canonical morphism
@@
 is an almost weak equivalence.
-  • proof.
-Let $L^{\rm al}:\mathcal{M}\to\mathcal{M}^{\rm al}$ be another Bousfield localization by almost weak equivalences.
-By the universality of $L^{\rm al}$, the natural transformation $\mathrm{Id}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ homotopically factors as
+[…]
 […]
 By (2) in Proposition 2.11, the composition $\psi(L^{\rm al})\circ\varphi(L^{\rm al}):L^{\rm al}\to L^{\rm al}\circ L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,L^{\rm al}(-))$ is an equivalence.
-Therefore, the induced transformation $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,\varphi):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,L^{\rm al}(-))$ is identified with a homotopy section of $\psi:L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$.
-Further, since the functor $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ is homotopically idempotent, the factorization
-$\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,\psi)\circ\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,\varphi):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,L^{\rm al}(-))\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\circ\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$
-implies that $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ is a homotopy retraction of $L^{\rm al}$.
-Hence, the natural transformation $\psi:L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ is a weak equivalence.
-We show that the functor $\tilde{I}\otimes Q:\mathcal{M}\to\mathcal{M}$ induced a left Quillen equivalence $\tilde{I}\otimes Q:\mathcal{M}^{\rm al}\to\mathcal{M}^{\rm fi}$.
-Since the functor $\tilde{I}\otimes Q:\mathcal{M}\to\mathcal{M}$ sends almost weak equivalences to weak equivalences, it factors through $\mathcal{M}^{\rm al}$.
-By (1) in Proposition 2.10 and the equivalence $\psi:L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$, the unit $\mathrm{Id}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\circ(\tilde{I}\otimes Q)$ induces an equivalence on the homotopy category of $\mathcal{M}^{\rm al}$, implying that it is a left Quillen equivalence.
-$\Box$
+Therefore, the induced transformation
+$\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,\varphi):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,L^{\rm al}(-))$
+[…]
+[…]
+[…]
+[…]
 3. Commutative monoid objects in almost mathematics of pointed closed symmetric monoidal model categories
@@
 Assume that $\tilde{\mathfrak{m}}$ is a flat $V$-module.
-In Smith ideal theory language, a $V$-homomorphism $\tilde{j}:\tilde{\mathfrak{m}}\to V$ is regarded a commutative Smith ideal of $V$.
-For any $V$-algebra $A$, the almost unitalization $A_{!!}$ can be represented by the pushout product
+In the language of Smith ideal theory, a $V$-homomorphism $\tilde{j}:\tilde{\mathfrak{m}}\to V$ is regarded as a commutative Smith ideal of $V$.
+For any $V$-algebra AA, the almost unitalization $A_{!!}$ can be represented by the pushout product
 $\tilde{j}\Box(V\to A)=(V\amalg_{\tilde{\mathfrak{m}}}({\tilde{\mathfrak{m}}\otimes A}))\to A,$
 which is an almost isomorphism.
-In Gabber–Ramero’s text [GR03], $A_{!!}$ is defined by to be the source of $\tilde{j}\Box(V\to A)$ without Smith ideal theory.
-On the view point of Smith ideal theory, we can easily check that the induced ring homomorphism $A_{!!}\to A$ is an almost isomorphism.
+In Gabber–Ramero’s text [5], $A_{!!}$ is defined as the source of $\tilde{j}\Box(V\to A)$ without using Smith ideal theory.
+From the viewpoint of Smith ideal theory, one can easily check that the induced ring homomorphism $A_{!!}\to A$ is an almost isomorphism.
 Indeed, by the associativity of pushout products,
-$(\tilde{\mathfrak{m}}\otimes_{V}{A}_{!!}\to\tilde{\mathfrak{m}}\otimes_{V}A)\simeq L_{1}(\tilde{\mathfrak{m}})\Box(\tilde{j}\Box(V\to A))\to(L_{1}(\tilde{\mathfrak{m}})\Box\tilde{j})\Box(V\to A)\simeq\mathrm{id}_{\tilde{\mathfrak{m}}}\Box(V\to A)\simeq\mathrm{id}_{\tilde{\mathfrak{m}}\otimes_{V}A}$
+$(\tilde{\mathfrak{m}}\otimes_{V}{A}_{!!}\to\tilde{\mathfrak{m}}\otimes_{V}A)\simeq L_{1}(\tilde{\mathfrak{m}})\Box(\tilde{j}\Box(V\to A))\\ \to(L_{1}(\tilde{\mathfrak{m}})\Box\tilde{j})\Box(V\to A)\simeq\mathrm{id}_{\tilde{\mathfrak{m}}}\Box(V\to A)\simeq\mathrm{id}_{\tilde{\mathfrak{m}}\otimes_{V}A}$
 is an isomorphism between arrows, implying that those sources $\tilde{\mathfrak{m}}\otimes_{V}A_{!!}$ and $\tilde{\mathfrak{m}}\otimes_{V}A$ are isomorphic.
@@
 3.1. Almost unitalization functors of commutative algebra objects of symmetric monoidal model categories
-In general the adjunction $\mathrm{Sym}:\mathcal{M}\rightleftarrows\mathrm{CAlg}(\mathcal{M}):U$ does not induces a model structure on $\mathrm{CAlg}(\mathcal{M})$ but it has the semi-model structure defined in [Spi01, Definition 2.3] or [Whi22, Definition 3.9].
-Semi-model structure only needs that trivial cofibrations with cofibrant domains have the left lifting property with respect to all fibrations, and only maps with cofibrant domains factor into a trivial cofibration followed by a fibration.
-If all objects of a semi-model category are cofibrant, it is just a model category (See [Whi22, Definition 3.9]).
-Left and right Quillen functors between semi-model categories are defined as the same as ones of model categories.
-For any class of morphisms $\mathcal{I}$, let $\mathrm{Cell}(\mathcal{I})$ denote the class of morphisms that is a transfinite composition of pushouts of the element of $\mathcal{I}$, and an object of $\mathrm{Cell}(\mathcal{I})$ is said to be a relative $\mathcal{I}$-cell complex.
-Let $\mathcal{M}$ is cofibrantly generated with a set of generating cofibrations $\mathcal{I}$ and a set of generating trivial cofibrations $\mathcal{J}$.
-By White [Whi22, Corollary 3.8], if the domains of $\mathcal{I}$ (resp.
-$\mathcal{J}$) is small relative to $\mathrm{Cell}(\mathcal{I}\otimes\mathcal{M})$ (resp.
-$\mathrm{Cell}(\mathcal{J}\otimes\mathcal{M})$), then the category of commutative $R$-algebras has a semi-model structure inherited by $\mathcal{M}$ for any commutative monoid object $R$.
+In general the adjunction $\mathrm{Sym}:\mathcal{M}\rightleftarrows\mathrm{CAlg}(\mathcal{M}):U$ does not induces a model structure on $\mathrm{CAlg}(\mathcal{M})$ but it has the semi-model structure defined in [17, Definition 2.3] or [19, Definition 3.9].
+A semi-model structure only requires that trivial cofibrations with cofibrant domains have the left lifting property with respect to all fibrations, and only maps with cofibrant domains factor into a trivial cofibration followed by a fibration.
+If all objects of a semi-model category are cofibrant, it is just a model category (see [19, Definition 3.9]).
+Left and right Quillen functors between semi-model categories are defined in the same way as those of model categories.
+For any class of morphisms $\mathcal{I}$, let $\mathrm{Cell}(\mathcal{I})$ denote the class of morphisms that are transfinite compositions of pushouts of elements of $\mathcal{I}$, and a morphism in $\mathrm{Cell}(\mathcal{I})$ is said to be a relative $\mathcal{I}$-cell complex.
+Let $\mathcal{M}$ be cofibrantly generated with a set of generating cofibrations $\mathcal{I}$ and a set of generating trivial cofibrations $\mathcal{J}$.
+By White [19, Corollary 3.8], if the domains of $\mathcal{I}$ (resp.
+$\mathcal{J}$) are small relative to $\mathrm{Cell}(\mathcal{I}\otimes\mathcal{M})$ (resp.
+$\mathrm{Cell}(\mathcal{J}\otimes\mathcal{M})$), then the category of commutative $R$-algebras has a semi-model structure inherited from $\mathcal{M}$ for any commutative monoid object $R$.
 We recall the definition of lax monoidal Quillen adjunctions:
-Definition 3.1 ([SS03] Definition 3.6).
-Let $F:\mathcal{M}\rightleftarrows\mathcal{N}:U$ be a Quillen adjunction of monoidal model categories. Then we say that $(F,\,U)$ is lax Quillen monoidal if it satisfies the following properties:
+Definition 3.1 ([16] Definition 3.6).
+Let $F:\mathcal{M}\rightleftarrows\mathcal{N}:U$ be a Quillen adjunction of monoidal model categories. Then we say that (F,U)(F,\,U) is lax Quillen monoidal if it satisfies the following properties:
   • (1)
@@
 is lax Quillen monoidal.
-  • proof.
-    Since the class almost weak equivalences between cofibrant objects is closed under tensor product, the localization $L^{\rm al}:\mathcal{M}\to\mathcal{M}^{\rm al}$ is lax monoidal.
+Since the class of almost weak equivalences between cofibrant objects is closed under tensor product, the localization $L^{\rm al}:\mathcal{M}\to\mathcal{M}^{\rm al}$ is lax monoidal.
 Furthermore, given any pair of cofibrant objects $M$ and $N$, the product $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ induces a weak equivalence
@@
 The mapping functor $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-):\mathcal{M}\to\mathcal{M}^{\rm al}$ is lax monoidal, and the Quillen equivalence $(\tilde{I}\otimes Q,\,\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,R(-)))$ induces on their commutative monoid objects
-$\tilde{I}\otimes Q:\mathrm{CAlg}(\mathcal{M}^{\rm al})\rightleftarrows\mathrm{CAlg}(\mathcal{M}^{\rm fi}):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$
+[…]
 between semi-model categories. $\Box$
@@
 is a monoid morphism.
-  • proof.
 Replacing $V$ and $A$ and considering $\mathrm{CAlg}(\mathcal{M})_{A/}$, it is sufficient to prove that
@@
 $A_{!!}=\mathrm{Ev}_{0}((\tilde{j}:\tilde{I}\to V)\Box(L_{1}(A):V\to A))=V\amalg_{\tilde{I}}(\tilde{I}\otimes A)).$
-  • proof.
 Assume that $0\to A$ is a cofibration.
@@
 $\Box$
-We verify that, for any commutative monoid object $A$, the internal-hom object $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)$ has a commutative monoid structure on $\mathcal{M}$ inherited by one of $A$ and the homotopy section $\Delta:\tilde{I}\to\tilde{I}\otimes\tilde{I}$ satisfying the coherent conditions in Definition 2.2.
+We verify that, for any commutative monoid object $A$, the internal-hom object $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)$ has a commutative monoid structure in $\mathcal{M}$ inherited from that of $A$ and the homotopy section $\Delta:\tilde{I}\to\tilde{I}\otimes\tilde{I}$ satisfying the coherent conditions in Definition 2.2.
 Let $\mathrm{ev}_{\tilde{I}}:\tilde{I}\otimes\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)\to A$ denote the evaluation morphism induced by the identity on $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)$.
@@
 $\mu_{\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)}:\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)\otimes\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)\to\mathrm{Map}_{\mathcal{M}}(\tilde{I}\otimes\tilde{I},\,A)\overset{\Delta^{*}}{\to}\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A),$
-satisfying the associativity and unitality due to Condition (3) in Definition 2.2.
-In the case that the monoidal structure on $\mathcal{M}$ is closed, both the symmetric monoidal structures on the allow category $\mathcal{M}$ are closed by Hovey [Hov14, Theorem 1.2].
-For any couple of morphisms $(f:X_{0}\to X_{1},\,g:Y_{0}\to Y_{1})$ in $\mathcal{M}$, the internal-hom objects are determined as follows: The internal-hom object $\mathrm{Map}_{\mathcal{M}}^{\otimes}(f,\,g)$ of the diagonal product monoidal structure is defined by
+satisfying associativity and unitality by Condition (3) in Definition 2.2.
+In the case that the monoidal structure on $\mathcal{M}$ is closed, both symmetric monoidal structures on the arrow category $\mathrm{Ar}(\mathcal{M})$ are closed by Hovey [9, Theorem 1.2].
+For any pair of morphisms $(f:X_{0}\to X_{1},\,g:Y_{0}\to Y_{1})$ in $\mathcal{M}$, the internal-hom objects are determined as follows: The internal-hom object $\mathrm{Map}_{\mathcal{M}}^{\otimes}(f,\,g)$ of the diagonal product monoidal structure is defined by
 $\mathrm{Map}^{\otimes}(f,\,g):\mathrm{Map}_{\mathcal{M}}(X_{0},\,Y_{0})\times_{\mathrm{Map}_{\mathcal{M}}(X_{0},\,Y_{1})}\mathrm{Map}_{\mathcal{M}}(X_{1},\,Y_{1})\to\mathrm{Map}_{\mathcal{M}}(X_{1},\,Y_{1})$
@@
 Theorem 3.6.
-Let $\mathcal{M}$ be a pointed closed symmetric monoidal model category with a monoidal unit object $V$ and homotopically idempotent commutative Smith ideal $j:I\to V$.
+Let $\mathcal{M}$ be a pointed closed symmetric monoidal model category, and $V$ a monoidal unit of $\mathcal{M}$ with a homotopically idempotent commutative Smith ideal $j:I\to V$.
 Assume that $\mathcal{M}$ is left proper and combinatorial.
 Then the Quillen adjunction
-$\tilde{j}\Box(-):\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}))\rightleftarrows\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M})):\mathrm{Map}^{\Box}(\tilde{j},\,-)$
+[…]
 induces a Quillen equivalence
-$\tilde{j}\Box(-):\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al}))\rightleftarrows\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al})):\mathrm{Map}^{\Box}(\tilde{j},\,-),$
+[…]
 which is equivalent to the trivial Quillen equivalence
 $L_{0}\circ\mathrm{Ev}_{1}:\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al}))\rightleftarrows\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al})):U_{1}\circ\mathrm{Ev}_{0}.$
-  • proof.
 Since the subcategory $\mathrm{CAlg}(\mathcal{M})$ is closed under homotopy limits, for any morphism $f:A\to B$ of commutative monoid objects, $\mathrm{Map}^{\Box}(\tilde{j},\,f):A\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)\times_{\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,B)}B$ is a morphism of monoid objects by Corollary 3.3.
 The functor $\mathrm{Map}^{\Box}(\tilde{j},-):\mathrm{Ar}(\mathcal{M})\to\mathrm{Ar}(\mathcal{M})$ can be restricted on $\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}))$.
-By Proposition 3.5, $j\Box\mathrm{Id}\to\mathrm{id}_{Ev_{1}(-)}=L_{0}(\mathrm{Ev}_{1}((-)))$ is a projective equivalence in $\mathrm{CAlg}(\mathcal{M}^{\rm al})$.
-Therefore, their Quillen right adjoints $\mathrm{Map}^{\Box}(\tilde{j},\,-)$ and $U_{1}\circ\mathrm{Ev}_{0}$ are weakly equivalent on $\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al}))$.
-$\Box$
-We obtain an analogue of Gabber–Ramero [GR03, p.22, Proposition 2.2.29] as avv corollary of Theorem 3.6.
+[…]
+We obtain an analogue of Gabber–Ramero [5, p.22, Proposition 2.2.29] as a corollary of Theorem 3.6.
 Corollary 3.7.
@@
 $\mathrm{Ev}_{0}(\tilde{j}\Box L_{1}(-)):\mathrm{CAlg}(\mathcal{M})\to\mathrm{CAlg}(\mathcal{M}).$
-  • proof.
-    Since all of $L_{1}$, $\tilde{j}\Box(-)$, and $\mathrm{Ev}_{0}$ are left Quillen functors, the composition $(-)_{!!}$ is also a left Quillen functor, and its right adjoint is the composition of their right adjoint functors
+[…]
 $\mathrm{Ev}_{1}\circ\mathrm{Map}^{\Box}(\tilde{j},\,-)\circ U_{0}:\mathrm{CAlg}(\mathcal{M})\to\mathrm{CAlg}(\mathcal{M}).$
@@
 In a stable model category, a homotopy Cartesian square is also homotopy coCartesian and vice versa.
-Further, in this section, we consider the case that stable monoidal model categories are generated by the monoidal unit objects.
-Those such categories said to be monogenic, and we explain in Definition 4.1.
-In this section, we prove that Bousfield bilocalization functors determine homotopically idempotent Smith ideals under the condition that those model categories are stable closed symmetric monoidal and monogenic.
-Furthermore, we show that homotopically closed objects with respect to those determined Smith ideals by Bousfield bilocalization are homotopically coincide with local objects.
+Further, in this section, we consider stable monoidal model categories generated by their monoidal unit objects.
+Such categories are called monogenic, as explained in Definition 4.1.
+In this section, we prove that Bousfield bilocalization functors determine homotopically idempotent Smith ideals under the condition that the model categories are stable, closed symmetric monoidal, and monogenic.
+Furthermore, we show that homotopically closed objects with respect to the Smith ideals determined by Bousfield bilocalization homotopically coincide with local objects.
 4.1. Monogenic model categories
-Following MacLane [Mac88, Chapter V, Section 7, p.127], we recall the definition of generators of categories: Let $\mathcal{M}$ be a category and $\mathcal{G}$ a class of small objects of $\mathcal{M}$.
+Following MacLane [11, Chapter V, Section 7, p.127], we recall the definition of generators of categories: Let $\mathcal{M}$ be a category and $\mathcal{G}$ a class of small objects of $\mathcal{M}$.
 The set $\mathcal{G}$ is a generator of $\mathcal{M}$ if for any pair of morphisms $h,\,h^{\prime}:X\to Y$ in $\mathcal{M}$, $h\neq h^{\prime}$ implies that there exists $S\in\mathcal{G}$ and a morphism $f:S\to X$ such that $h\circ f\neq h^{\prime}\circ f$.
-As Hovey’s book [Hov99, p.183, Definition 7.2.1], we define weakly generators of model categories.
-Definition 4.1 (c.f. [HPS97] p.7).
+Following Hovey’s book [8, p.183, Definition 7.2.1], we define weak generators of model categories.
+Definition 4.1 (c.f. Hovey [8]).
 Let $\mathcal{M}$ be a model category with a class $\mathcal{G}$ of objects of $\mathcal{M}$.
@@
 A symmetric monoidal model category weakly generated by the small monoidal unit is said to be monogenic.
-By the following proposition, for stable model categories, the above definition coincides with [Hov99, p.183, Definition 7.2.1].
+By the following proposition, for stable model categories, the above definition coincides with [8, p.183, Definition 7.2.1].
 Proposition 4.2.
@@
 Then an object $M$ is contractible if and only if the internal hom-object $\mathrm{Map}_{\mathcal{M}}(S,\,M)$ is contractible for any $S\in\mathcal{G}$.
-  • proof.
 The if-direction is clear.
 Let $M$ be an object of $\mathcal{M}$ such that the internal-hom object $\mathrm{Map}_{\mathcal{M}}(S,\,M)$ is contractible for any $S\in\mathcal{G}$.
-Since $\mathcal{G}$ generates $\mathcal{M}$, for any object $N$ and morphism $f:N\to M$, the condition that $f$ is not null-homotopic implies that there exists an object $S\in\mathcal{G}$ and a morphism $i:S\to N$ and $f\circ i:S\to M$ is not null-homotopic, contradicting the generating condition.
+Since $\mathcal{G}$ generates $\mathcal{M}$, for any object $N$ and morphism $f:N\to M$, the condition that $f$ is not null-homotopic implies that there is an object $S\in\mathcal{G}$ and a morphism $i:S\to N$ such that $f\circ i:S\to M$ is not null-homotopic, contradicting the generating condition.
 $\Box$
@@
 A functor $F:\mathcal{M}\to\mathcal{N}$ is said to be a Bousfield bilocalization if it is a Bousfield localization admitting fully faithful left adjoint and fully faithful right adjoint such that all of those functors are Quillen functors.
-We consider a Bousfield bilocalization $F:\mathcal{M}\to L(\mathcal{M})$ and let denote $F_{*}$ its fully faithful left adjoint and $U$ its fully faithfully right
-${F:\mathcal{M}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces L(\mathcal{M}).}$$\scriptstyle{U}$$\scriptstyle{F_{*}}$
+We consider a Bousfield bilocalization $F:\mathcal{M}\to L(\mathcal{M})$ and let $F_{*}$ denote its fully faithful left adjoint and $U$ its fully faithful right adjoint, respectively.
+[…]
 Further, we assume that the Quillen adjunction
@@
 $F_{*}:L(\mathcal{M})\rightleftarrows\mathcal{M}:F$
-is lax Quillen monoidal. Let $j:I\to V$ be a homotopy image object of $q(V)\circ c_{Q(V)}:F_{*}(F(Q(V))\to Q(V)\to V$. Then all of the canonical morphisms in the diagram
-$I\otimes I\rightarrow F_{*}(F(Q(V))\otimes I\rightarrow F_{*}(F(Q(V))\otimes F_{*}(F(Q(V))\rightarrow F_{*}(F(Q(V)))\leftarrow I$
+is lax Quillen monoidal. Let $j:I\to V$ be a homotopy image object of $q(V)\circ c_{Q(V)}:F_{*}(F(Q(V)))\to Q(V)\to V$. Then all of the canonical morphisms in the diagram
+$I\otimes I\rightarrow F_{*}(F(Q(V)))\otimes I\rightarrow F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))\rightarrow F_{*}(F(Q(V)))\leftarrow I$
 are weak equivalences.
-  • proof.
 Let $q(V):Q(V)\to V$ be a cofibrant replacement of $V$.
-Since the Bousfield localization $F$ is a left and right Quillen functor, $F(Q(V))\to F(V)$ is a trivial fibration, and $F(Q(V))$ is a cofibrant object of $L(\mathcal{M})$.
+Since the Bousfield localization $F$ is both a left and a right Quillen functor, $F(Q(V))\to F(V)$ is a trivial fibration, and $F(Q(V))$ is a cofibrant object of $L(\mathcal{M})$.
 Furthermore, since $F_{*}$ is lax Quillen monoidal, $F(Q(V))$ is equivalent to a monoidal unit of $L(\mathcal{M})$.
@@
 $F(Q(V))\otimes F(Q(V))\to F(V)\otimes F(Q(V))\to F(Q(V))\to F(V)$
-    are weak equivalences by Ken Brown’s lemma [Hov99, p.6 Lemma 1.1.12]. In particular,
+are weak equivalences by Ken Brown’s lemma [8, p.6 Lemma 1.1.12]. In particular,
 $\mu:F(Q(V))\otimes F(Q(V))\to F(Q(V))$
-    is a weak equivalence between cofibrant objects. Since $F_{*}$ is a left Quillen functor, by Ken Brown’s lemma [Hov99, p.6 Lemma 1.1.12] again,
+is a weak equivalence between cofibrant objects. Since $F_{*}$ is a left Quillen functor, by Ken Brown’s lemma [8, p.6 Lemma 1.1.12] again,
 $F_{*}(\mu):F_{*}(F(Q(V))\otimes F(Q(V)))\to F_{*}(F(Q(V)))$
@@
 $\mu_{F_{*}(F(Q(V)))}:F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))\to F_{*}(F(Q(V)))$
-    is homotopic to the pushout of $\nabla:F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V))\to F_{*}(F(Q(V))\otimes F(Q(V)))$ along the morphism $F_{*}(F(\mu)):F_{*}(F(Q(V))\otimes F(Q(V)))\to F_{*}(F(Q(V)))$.
-By the left properness of $\mathcal{M}$ implies that $\mu:F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))\to F_{*}(F(Q(V)))$ is a weak equivalence.
-By the stability of $\mathcal{M}$, the canonical morphism $(j:I\to V)\to(q(V)\circ c_{Q(V)}:F_{*}(F(Q(V)))\to V)$ is a weak equivalence in the projective model structure.
-Since $I\to F_{*}(F(Q(V)))$ is a weak equivalence of cofibrant objects, by Ken Brown’s lemma again, both morphisms in $I\otimes I\rightarrow F_{*}(F(Q(V))\otimes I\rightarrow F_{*}(F(Q(V))\otimes F_{*}(F(Q(V))$ are weak equivalences.
+is homotopic to the pushout of $\nabla:F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))\to F_{*}(F(Q(V))\otimes F(Q(V)))$ along the morphism $F_{*}(F(\mu)):F_{*}(F(Q(V))\otimes F(Q(V)))\to F_{*}(F(Q(V)))$.
+By the left properness of $\mathcal{M}$, $\mu:F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))\to F_{*}(F(Q(V)))$ is a weak equivalence.
+By the stability of $\mathcal{M}$, the morphism $(j:I\to V)\to(q(V)\circ c_{Q(V)}:F_{*}(F(Q(V)))\to V)$ is a weak equivalence in the projective model structure.
+Since $I\to F_{*}(F(Q(V)))$ is a weak equivalence of cofibrant objects, by Ken Brown’s lemma again, both morphisms in $I\otimes I\rightarrow F_{*}(F(Q(V)))\otimes I\rightarrow F_{*}(F(Q(V)))\otimes F_{*}(F(Q(V)))$ are weak equivalences.
 $\Box$
 Proposition 4.5.
-Let $F:\mathcal{M}\to L(\mathcal{M})$ be a Bousfield bilocalization of stable symmetric monoidal model category, and $F_{*}$ denote its fully faithful left adjoint. Assume that the Quillen adjunction
+Let $F:\mathcal{M}\to L(\mathcal{M})$ be a Bousfield bilocalization of a stable symmetric monoidal model category, and $F_{*}$ denote its fully faithful left adjoint.
+Assume that the Quillen adjunction
 $F_{*}:L(\mathcal{M})\rightleftarrows\mathcal{M}:F$
 is lax Quillen monoidal.
-Then the counit $c:F_{*}\circ F\to\mathrm{Id}_{\mathcal{M}}$ induces determines a homotopically idempotent commutative Smith ideal $j:I\to V$ of a monoidal unit object $V$, where $I$ is a cofibrant replacement of the homotopy image of $q(V)\circ c_{Q(V)}:F_{*}(F(Q(V))\to Q(V)\to V$.
-  • proof.
+Then the counit $c:F_{*}\circ F\to\mathrm{Id}_{\mathcal{M}}$ determines a homotopically idempotent commutative Smith ideal $j:I\to V$ of a monoidal unit object $V$, where $I$ is a cofibrant replacement of the homotopy image of $q(V)\circ c_{Q(V)}:F_{*}(F(Q(V)))\to Q(V)\to V$.
 Since $\mathcal{M}$ is stable, the unit morphism $c_{V}\to j$ is a projective weak equivalence in the arrow category of $\mathrm{Ar}(\mathcal{M})$.
@@
 Proposition 4.6.
-Let $\mathcal{M}$ be a pointed closed symmetric monoidal model with a monoidal unit $V$ and $F:\mathcal{M}\to\mathcal{M}$ a symmetric monoidal Bousfield bilocalization, and $F_{*}:\mathcal{M}\to\mathcal{M}$ denote the homotopically fully faithful left adjoint.
+Let $\mathcal{M}$ be a pointed closed symmetric monoidal model category with a monoidal unit $V$ and $F:\mathcal{M}\to L(\mathcal{M})$ a symmetric monoidal Bousfield bilocalization, and $F_{*}:L(\mathcal{M})\to\mathcal{M}$ denote the homotopically fully faithful left adjoint.
 Assume that the Quillen adjunction
@@
 is a weak equivalence for any object $M$.
-  • proof.
 Let $j:I\to F_{*}(F(V))$ be a cofibrant replacement of the homotopy image of the counit $c_{V}:F_{*}(F(V))\to V$.
@@
 Theorem 4.7.
-Let $\mathcal{M}$ be a stable closed symmetric monoidal model category and $F:\mathcal{M}\to L(\mathcal{M})$ a Bousfield bilocalization, and $F_{*}:\mathcal{M}\to\mathcal{M}$ denote the homotopically fully faithful left adjoint.
+Let $\mathcal{M}$ be a stable closed symmetric monoidal model category and $F:\mathcal{M}\to L(\mathcal{M})$ a Bousfield bilocalization, and $F_{*}:L(\mathcal{M})\to\mathcal{M}$ denote the homotopically fully faithful left adjoint.
 Assume that $\mathcal{M}$ is monogenic and the Quillen adjunction
@@
 is lax Quillen monoidal.
-Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal determined as a result of Proposition 4.5.
+Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal obtained in Proposition 4.5.
 Let $U$ denote a right Quillen adjoint of $F$.
-Then all of the morphisms of the diagram
-[…]
+Then all morphisms in the diagram
+$UF(M)\overset{\varphi}{\rightarrow}UF\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,M)\overset{\psi}{\leftarrow}\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,M)$
 are weak equivalences for any fibrant object $M$ of $\mathcal{M}$.
-  • proof.
-    Let $j:\tilde{I}\to V$. For any fibrant object $M$ of $\mathcal{M}$, one has a commutative diagram
-[…]
+Let $j:I\to V$. For any fibrant object $M$ of $\mathcal{M}$, one has a commutative diagram
+${\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(V,\,UF(M))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\hskip-14.22636pt\varphi_{*}}$${\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(V,\,UF\mathrm{Map}_{\mathcal{M}}(I,\,M))}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(V,\,\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,M))}$$\scriptstyle{\hskip 14.22636pt\psi_{*}}$${\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(F_{*}(F(Q(V))),\,M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(F_{*}(F(Q(V)))\otimes I,\,M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\mathrm{Hom}_{\mathrm{Ho}(\mathcal{M})}(\tilde{I},\,M),}$
 where all vertical maps are isomorphisms.
@@
 Remark 5.1.
-The definitions of colors, symmetric sequences, and colored operads in symmetric monoidal categories can be found in Yau’s book, specifically in [Yau16, Chapter 9 and 11].
-Colored operads and maps colored operads are defined in [Yau16, p.176, Definition 11.2.1] and [Yau16, p.180, Definition 11.2.12], respectively.
-Additionally, the material of algebras over colored operads is explained in [Yau16, p.217, Definition 13.2.1].
-Further, in the case that the monoidal structure on $\mathcal{M}$ is closed, algebras over $\mathcal{O}$ are determined by operad maps from $\mathcal{O}$ to endomorphism operads [Yau16, p.226, Definition 13.8.1].
+The definitions of colors, symmetric sequences, and colored operads in symmetric monoidal categories can be found in standard references on operads; see, for example, White–Yau [18, Section 3].
+Further, in the case that the monoidal structure on $\mathcal{M}$ is closed, algebras over $\mathcal{O}$ are determined by operad maps from $\mathcal{O}$ to endomorphism operads.
 For any $\mathfrak{C}$-colored object $A=\{A_{c}\}_{c\in\mathfrak{C}}$ of $\mathcal{M}$, the endomorphism operad $\mathbf{End}_{\mathcal{M}}(A)$ is defined by
 $\mathbf{End}_{\mathcal{M}}(A)(c_{1},\ldots,c_{n};d)=\mathrm{Map}_{\mathcal{M}}(A_{c_{1}}\otimes\cdots\otimes A_{c_{n}},A_{d}).$
-Giving a map $\theta:\mathcal{O}\to\mathbf{End}_{\mathcal{M}}(A)$ of $\mathfrak{C}$-colored operads determines an $\mathcal{O}$-algebra structure on $A$. (See [Yau16, p.231, Proposition 13.9.1]).
+Giving a map $\theta:\mathcal{O}\to\mathbf{End}_{\mathcal{M}}(A)$ of $\mathfrak{C}$-colored operads determines an $\mathcal{O}$-algebra structure on $A$.
 Definition 5.2.
@@
 Since $L_{1}$ (resp.
-$L_{0}$) is strict monoidal (See [WY24, Definition 3.1.1 (7)]), $\overrightarrow{\mathcal{O}^{\Box}}$ (resp.
+$L_{0}$) is strict monoidal (See [18, Definition 3.1.1 (7)]), $\overrightarrow{\mathcal{O}^{\Box}}$ (resp.
 $\overrightarrow{\mathcal{O}^{\otimes}}$) is a colored operad in $\mathrm{Ar}^{\Box}(\mathcal{M})$ (resp.
@@
 $\mathcal{O}^{\otimes}$-algebra objects of $\mathrm{Ar}^{\otimes}(\mathcal{M})$).
-Operadic Smith ideals defined as the following:
-Definition 5.3 ([WY24] Definition 3.3.1).
+Operadic Smith ideals are defined as follows:
+Definition 5.3 ([18] Definition 3.3.1).
 Let $\mathcal{M}$ be a symmetric monoidal closed (model) category and $\mathcal{O}$ a colored operad in $\mathcal{M}$.
@@
 5.2. Non-unital algebras from the (operadic) Smith ideal theory viewpoint
-We recall a brief review of non-unital rings, which are rings without assuming the existence of the multiplicative unit $1$.
+We briefly review non-unital rings, which are rings not assumed to have a multiplicative unit $1$.
 Any unital ring is regarded as a non-unital ring by forgetting the existence of $1$.
-An augmentation $\varepsilon_{A}:A\to V$ is a ring homomorphism that the unit homomorphism $V\to A$ is a section of $\varepsilon_{A}$, and a $V$-algebra with an augmentation is called an augmented $V$-algebra.
-Let $\mathrm{Alg}_{V//V}$ denotes the category of augmented $V$-algebras and $\mathrm{Alg}^{\rm nu}_{V}$ the category of non-unital $V$-algebras.
-It is well-known that there exists a categorical equivalence between $\mathrm{Alg}^{\rm nu}_{V}$ and $\mathrm{Alg}_{V//V}$ as follows: For any $V$-algebra $A$, the direct sum $V\oplus A$ has a canonical unital ring structure defined by
+An augmentation $\varepsilon_{A}:A\to V$ is a ring homomorphism such that the unit homomorphism $V\to A$ is a section of $\varepsilon_{A}$, and a $V$-algebra with an augmentation is called an augmented $V$-algebra.
+Let $\mathrm{Alg}_{V//V}$ denote the category of augmented $V$-algebras and $\mathrm{Alg}^{\rm nu}_{V}$ the category of non-unital $V$-algebras.
+It is well-known that there exists a categorical equivalence between $\mathrm{Alg}^{\rm nu}_{V}$ and $\mathrm{Alg}_{V//V}$ as follows: For any $V$-algebra AA, the direct sum $V\oplus A$ has a canonical unital ring structure defined by
 $(m,\,a)\cdot(n,\,b)=(mn,\,na+mb+ab)$
-for $m,\,n\in V$ and $a,\,b\in A$. It is easily checked that the functor the adjunction
+for $m,\,n\in V$ and $a,\,b\in A$. It is easily checked that the adjunction
 $V\oplus(-):\mathrm{Alg}^{\rm nu}_{V}\rightleftarrows\mathrm{Alg}_{V//V}:\mathrm{Ker}(\epsilon_{(-)}:-\to V)$
@@
 From the Smith ideal theory viewpoint, the above categorical equivalence is a special case of cokernel-and-kernel correspondence between the arrow categories.
-We consider the case abelian categories.
-Let $\mathcal{M}$ be a symmetric closed monoidal abelian category and $\mathrm{Ar}^{\rm im}(\mathcal{M})$ denote the full subcategory spanned those objects $f:R\to S$ such that the unit $f\to\mathrm{ker}(\mathrm{cok}(f))$ is an isomorphism.
+We consider the case of abelian categories.
+Let $\mathcal{M}$ be a symmetric closed monoidal abelian category and $\mathrm{Ar}^{\rm im}(\mathcal{M})$ denote the full subcategory spanned by those objects $f:R\to S$ such that the unit $f\to\mathrm{ker}(\mathrm{cok}(f))$ is an isomorphism.
 Then the cokernel functor $\mathrm{Ar}^{\rm im}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$ is a categorical equivalence.
@@
 Let $\mathcal{M}$ be a locally presentable abelian symmetric monoidal category.
-Then, the arrow category $\mathrm{Ar}(\mathcal{M})$ is also locally presentable and there exist reflective localization functors $L_{\rm im}:\mathrm{Ar}^{\Box}(\mathcal{M})\to\mathrm{Ar}^{\rm im}(\mathcal{M})$ by the unit $\mathrm{id}\to\ker\circ\mathrm{coker}$ and $L_{\rm coim}:\mathrm{Ar}^{\otimes}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$ by the counit $\mathrm{coker}\circ\ker\to\mathrm{id}$ such that the adjunction
+Then the arrow category $\mathrm{Ar}(\mathcal{M})$ is also locally presentable and there exist reflective localization functors $L_{\rm im}:\mathrm{Ar}^{\Box}(\mathcal{M})\to\mathrm{Ar}^{\rm im}(\mathcal{M})$ by the unit $\mathrm{id}\to\ker\circ\mathrm{coker}$ and $L_{\rm coim}:\mathrm{Ar}^{\otimes}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$ by the counit $\mathrm{coker}\circ\ker\to\mathrm{id}$ such that the adjunction
 $\mathrm{cok}:\mathrm{Ar}^{\Box}(\mathcal{M})\rightleftarrows\mathrm{Ar}^{\otimes}(\mathcal{M}):\mathrm{ker}$
@@
 $\mathrm{cok}:\mathrm{Ar}^{\rm im}(\mathcal{M})\rightleftarrows\mathrm{Ar}^{\rm coim}(\mathcal{M}):\mathrm{ker}.$
-  • proof.
 Since any abelian category is binormal, $f\to\mathrm{im}(f)$ is an isomorphism if and only if $f$ is a monomorphism.
-By Adámek–Rosický [AR94, p.44, Corollary 1.5.4], the arrow category is locally presentable, and it admits reflective localization.
-By the definitions of those functors: $\mathrm{im}=\mathrm{ker}\circ\mathrm{cok}$ and $\mathrm{coim}=\mathrm{cok}\circ\mathrm{ker}$, the restriction $\mathrm{cok}:\mathrm{Ar}^{\rm im}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$ are quasi-inverse functors.
+By Adámek–Rosický [1, p.44, Corollary 1.5.4], the arrow category is locally presentable, and it admits reflective localization.
+By the definitions of those functors: $\mathrm{im}=\mathrm{ker}\circ\mathrm{cok}$ and $\mathrm{coim}=\mathrm{cok}\circ\mathrm{ker}$, the restrictions of $\mathrm{cok}$ and $\mathrm{ker}$ to $\mathrm{Ar}^{\rm im}(\mathcal{M})$ and $\mathrm{Ar}^{\rm coim}(\mathcal{M})$ are quasi-inverse functors.
 $\Box$
@@
 Let $\mathcal{M}$ be a locally presentable symmetric monoidal abelian category with a monoidal unit $V$.
-A commutative non-unital monoid object of $\mathcal{M}$ is a commutative Smith ideal $j:I\to R$ in the localized full subcategory $\mathrm{Ar}^{\rm im}(\mathcal{M})$ satisfying that the cokernel $\mathrm{cok}(j):R\to\mathrm{Coker}(j)$ is isomorphic with an argumentation $\varepsilon_{R}:R\to V$, and $\mathrm{Alg}_{\mathcal{O}}^{\rm nu}(\mathcal{M})$ denote the full subcategory of $\mathrm{Alg}_{\mathcal{O}^{\Box}}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ spanned by non-unital $\mathcal{O}^{\Box}$-algebra objects.
+A commutative non-unital monoid object of $\mathcal{M}$ is a commutative Smith ideal $j:I\to R$ in the localized full subcategory $\mathrm{Ar}^{\rm im}(\mathcal{M})$ satisfying that the cokernel $\mathrm{cok}(j):R\to\mathrm{Coker}(j)$ is isomorphic to an augmentation $\varepsilon_{R}:R\to V$, and $\mathrm{Alg}_{\mathcal{O}}^{\rm nu}(\mathcal{M})$ denotes the full subcategory of $\mathrm{Alg}_{\mathcal{O}^{\Box}}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ spanned by non-unital $\mathcal{O}^{\Box}$-algebra objects.
 In particular, $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ denotes the full subcategory $\mathrm{CAlg}(\mathrm{Ar}^{\Box}(\mathcal{M}))$ spanned by commutative Smith ideals whose cokernels are isomorphic to $V$.
@@
 Proposition 5.6.
-Let $\mathcal{M}$ be a locally presentable symmetric monoidal abelian category with a monoidal unit $V$. The categorical equivalences
+Let $\mathcal{M}$ be a locally presentable symmetric monoidal abelian category with a monoidal unit $V$. The categorical equivalence
 $\mathrm{cok}:\mathrm{Ar}^{\rm im}(\mathcal{M})\to\mathrm{Ar}^{\rm coim}(\mathcal{M})$
@@
 $\mathrm{Alg}^{\rm nu}_{\mathcal{O}^{\Box}}(\mathcal{M})\to\mathrm{Alg}_{\mathcal{O}^{\otimes}}(\mathcal{M})_{V//V}.$
-  • proof.
 By definition of the category $\mathrm{Alg}_{\mathcal{O}^{\Box}}^{\rm nu}(\mathcal{M})$, one has an isomorphism $V\oplus I\to R$. The assertion is clear. $\Box$
@@
 Let $\mathcal{M}$ be a symmetric monoidal closed model category with a monoidal unit $V$.
-An homotopically augmented $\mathcal{O}$-algebra over $V$ $f:A\to B$ is an object of $\mathrm{Alg}_{\overrightarrow{\mathcal{O}^{\otimes}}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))$ that the unit $U^{0}(V)=(\mathrm{id}_{V}:V\to V)\to(f:A\to B)$ has a homotopy section.
-Let $\mathrm{Alg}_{\mathcal{O}^{\otimes}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))_{V//V}$ denotes the semi-model category of homotopically augmented $V$-algebras.
+An homotopically augmented $\mathcal{O}$-algebra over $V$ $f:A\to B$ is an object of $\mathrm{Alg}_{\overrightarrow{\mathcal{O}^{\otimes}}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))$ such that the unit $U^{0}(V)=(\mathrm{id}_{V}:V\to V)\to(f:A\to B)$ has a homotopy section.
+Let $\mathrm{Alg}_{\mathcal{O}^{\otimes}}(\mathrm{Ar}^{\otimes}(\mathcal{M}))_{V//V}$ denote the semi-model category of homotopically augmented $V$-algebras.
 In particular, $\mathrm{CAlg}(\mathcal{M})_{V//V}$ denotes the semi-model category of homotopically augmented $V$-algebras.
@@
 In particular, $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ denotes the semi-model category of non-unital commutative $V$-algebra object of $\mathcal{M}$.
-Due to the following White and Yau’s theorem, we have a Quillen equivalence between semi-model categories of non-unital non-unital $\mathcal{O}$-algebras and homotopically augmented $\mathcal{O}$-algebras.
-Theorem 5.8 ([WY24] Theorem 4.4.1).
+By the following theorem of White and Yau, we have a Quillen equivalence between semi-model categories of non-unital $\mathcal{O}$-algebras and homotopically augmented $\mathcal{O}$-algebras.
+Theorem 5.8 ([18] Theorem 4.4.1).
 Let $\mathcal{M}$ be a stable symmetric monoidal model category with a monoidal unit $V$.
@@
 White introduced the commutative monoid axiom of monoidal model categories.
-Definition 5.10 ([Whi22], Definition 3.1 and 3.4).
-A symmetric monoidal model category $\mathcal{M}$ is said to be satisfying the commutative monoid axiom if, for all $n>0$, the $n$-fold symmetric pushout product functor $(-)^{\Box n}/\Sigma_{n}:\mathcal{M}\to\mathcal{M}$ preserves all trivial cofibrations.
-Further if the $n$-fold symmetric pushout product functor preserves all cofibrations and trivial cofibrations, $\mathcal{M}$ is said to be satisfying the strong commutative monoid axiom.
-In particular, by [WY24, Corollary 5.1.2], we obtain the following:
+Definition 5.10 ([19], Definition 3.1 and 3.4).
+A symmetric monoidal model category $\mathcal{M}$ is said to satisfy the commutative monoid axiom if, for all $n>0$, the $n$-fold symmetric pushout product functor $(-)^{\Box n}/\Sigma_{n}:\mathcal{M}\to\mathcal{M}$ preserves all trivial cofibrations.
+Further, if the $n$-fold symmetric pushout product functor preserves all cofibrations and trivial cofibrations, $\mathcal{M}$ is said to satisfy the strong commutative monoid axiom.
+In particular, by [18, Corollary 5.1.2], we obtain the following:
 Corollary 5.11.
-Let $\mathcal{M}$ be a cofibrantly generated stable model category satisfying the strong commutative monoid axiom, the monoid axiom, and in which cofibrant commutative $\mathrm{Ar}^{\Box}(\mathcal{M})$.
+Let $\mathcal{M}$ be a cofibrantly generated stable model category satisfying the strong commutative monoid axiom and the monoid axiom, and assume that the category of commutative monoids in $\mathrm{Ar}^{\Box}(\mathcal{M})$ has the transferred model structure.
 Then the cokernel-and-kernel adjunction induces a Quillen equivalence
@@
 Remark 5.12.
-If $\mathcal{M}$ is a locally presentable stable symmetric monoidal model category, the presentable $\infty$-category of augmented algebra is represented by the model category $\mathrm{CAlg}(\mathcal{M})_{V//V}$.
-Hence the left-hand-side $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ is equivalent to the $\infty$-category defined by Lurie [Lur17, p.949, Definition 5.4.4.9 and Proposition 5.4.4.10].
+If $\mathcal{M}$ is a locally presentable stable symmetric monoidal model category, the presentable $\infty$-category of augmented algebra is represented by $\mathrm{CAlg}(\mathcal{M})_{V//V}$.
+Hence the left-hand side $\mathrm{CAlg}^{\rm nu}(\mathcal{M})$ is equivalent to the $\infty$-category defined by Lurie [10, p.949, Definition 5.4.4.9 and Proposition 5.4.4.10].
+Acknowledgements
+I would like to thank anonymous referees for giving me innovative ideas and constructive comments to use Hovey’s Smith ideal theory and White and Yau’s operadic Smith ideal theory.
+I am also grateful to Isamu Iwanari for pointing out the derived-completion example.
+Use of AI tools.
+The author used Google’s Gemini-pro 3.1 as a conversational research aid for brainstorming and refining mathematical formulations, and OpenAI’s Prism (GPT-5.2) for editorial and expository assistance, including structural refinement.
+These tools were used to improve the manuscript’s clarity and to indicate where abbreviated arguments warranted fuller exposition.
+All results, proofs, and final formulations were reviewed, verified, and approved by the author, who bears sole responsibility for the content of this work.